### v1

# Non-normal Dynamics on Non-reciprocal Networks: Reactivity and Effective Dimensionality in Neural Circuits

Abstract
Non-reciprocal interactions are a defining feature of many complex systems, biological, ecological, and technological, often pushing them far from equilibrium and enabling rich dynamical responses. These asymmetries can arise at multiple levels: locally, in the dynamics of individual units, and globally, in the topology of their interactions. In this work, we investigate how these two forms of non-reciprocity interact in networks of neuronal populations. At the local level, each population is modeled by a non-reciprocally coupled set of excitatory and inhibitory neural populations exhibiting transient amplification and reactivity. At the network level, these populations are coupled via directed, asymmetric connections that introduce structural non-normality. Since non-reciprocal interactions generically lead to non-normal linear operators, we frame both local and global asymmetries in terms of non-normal dynamics. Using a modified Wilson–Cowan framework, we analyze how the interplay between these two types of non-normality shapes the system’s behavior. We show that their combination leads to emergent collective dynamics, including fluctuation-driven transitions, dimensionality reduction, and novel nonequilibrium steady states. Our results provide a minimal yet flexible framework to understand how multi-scale non-reciprocities govern complex dynamics in neural and other interconnected systems.
I Introduction
Non-reciprocity has recently emerged as a powerful organizing principle in the study of complex systems far from equilibrium. It describes the asymmetric, directed nature of interactions between components, whereby the influence of unit A on unit B is not necessarily matched by the reverse influence of B on A. This asymmetry underlies many forms of rich and often unexpected behavior across a broad range of domains, from active matter and synthetic meta-materials to neural circuits and ecological webs [8, 2, 9]. Non-reciprocal interactions break detailed balance and time-reversal symmetry at the microscopic level, enabling new types of collective dynamics — including, among others, spontaneous oscillations, traveling waves, and fluctuation-amplified responses — that do not exist in symmetric, equilibrium-like systems [4, 5, 10, 11].
In this paper, we aim to bring together and clarify two distinct, but often coexisting, manifestations of non-reciprocity in complex systems, both of which can be naturally interpreted through the lens of non-normality: a mathematical property of matrices where eigenvectors are non-orthogonal, allowing for transient amplification of perturbations. While non-normality provides a technical language to describe these effects, we emphasize that it is the underlying physical non-reciprocity — whether in the local dynamics or the network architecture — that drives the phenomena of interest.
The first form we address is dynamical non-normality, which arises at the level of the internal dynamics of individual units or subsystems. When the local Jacobian or stability matrix is non-normal, even stable fixed points can exhibit large transient responses to small perturbations, especially in the presence of noise [12, 13]. In neural circuits, for instance, populations of excitatory and inhibitory neurons often operate near critical balance, leading to strongly non-normal dynamics that transiently amplify fluctuations in input, contributing to fast signal processing, selectivity, and flexible population coding [14, 15, 16]. These effects are intrinsically linked to asymmetries in the internal coupling and can be viewed as a form of microscopic non-reciprocity between excitatory and inhibitory neurons.
The second form is what we term structural non-normality, which stems from the network-level asymmetry in the way local populations are connected at a larger scale. In many real-world systems, especially biological and engineered networks, the connectivity matrix that governs inter-unit influences is inherently non-reciprocal. Neural circuits, for example, are not fully bidirectional: higher-order cortical areas send feedback to early sensory regions, but not necessarily with equal weight or latency [17, 18, 19]. Similarly, ecological and social systems often exhibit trophic or influence hierarchies, where some magnitude flows in preferred directions [20]. When such asymmetries are strong, the overall network becomes structurally non-normal, leading to directional instabilities, long-range correlations, waves, and flow-like dynamics.
The central hypothesis we aim to explore is whether the interplay between local dynamical non-normality and global structural non-normality gives rise to qualitatively new collective behaviors, beyond what either form could produce in isolation. In other words, when locally non-reciprocal units (e.g., excitatory/inhibitory local neural populations) are embedded in a globally non-reciprocal network (e.g., feedforward/feedback cortical circuits), the system exhibits emergent properties that are not captured by standard linear response theory or symmetric interaction models. These might include enhanced stochastic amplification, novel transient modes, and selective propagation of fluctuations and signals across the network.
As a test of this hypothesis, here we analyze a minimal yet general neural model consisting of multiple interacting dynamical units, each described by a non-normal local system, coupled through a network, that in general is non-reciprocal. Using tools from linear stability analysis, stochastic dynamics, and spectral theory, we explore how the composition of non-reciprocities across scales governs the system’s susceptibility to noise, its capacity to transmit signals, and its repertoire of emergent dynamical states.
While our focus is primarily on neural systems —where both forms of non-reciprocity are particularly notorious— the conceptual framework we propose is broadly applicable to other domains, including active materials, biological networks, and collective decision-making systems. In all these contexts, non-reciprocity serves as a dynamical and structural engine for complexity, and understanding its layered manifestations is key to unlocking the behavior of high-dimensional, complex systems.
A theoretical neuroscience perspective:
II Wilson-Cowan model for neural populations
Dynamical non-normal effects at a single region
We consider the classical Wilson-Cowan model [6], which describes the dynamics of a large neural population composed of homogeneously interconnected excitatory and inhibitory neurons [21, 22, 13, 7]. The system of differential equations governing the dynamical evolution of the model is given by:
[…]
where $x(t)$ and $y(t)$ are the activation levels of the excitatory and inhibitory neuronal populations, respectively. The parameter $\alpha$ represents the rate of spontaneous activity decay, while $s$ denotes the net incoming current, $s(t)=\gamma_{\mu}x(t)-\gamma_{\nu}y(t)+h$, which is simply the sum of all synaptic excitatory and inhibitory inputs, weighted by their respective synaptic strengths —-$(\gamma_{\mu}$ for excitation and $\gamma_{\nu}$ for inhibition— plus a constant external input current, $h$. To reduce the dimensionality of the parameter space, all synaptic weights of a given type have been set to the same value. Finally $f(s)$ is a sigmoid non-negative response function. The system displays two distinct dynamical regimes: a quiescent (inactive) phase, where $\bar{x}=\bar{y}=0$, and an active phase $\bar{x}=\bar{y}>0$, characterized by nonzero steady-state activity with excitation dominating the response ($\bar{x},\bar{y}\neq 0$). These phases are separated by a bifurcation that occurs at the critical point $\gamma_{\mu}=\gamma_{\nu}+\alpha$: below this threshold, the system remains in the quiescent state, while above it, activity becomes self-sustained (see below).
In the deterministic limit, using variables $\Sigma=(x+y)/2$ and $\Delta=(x-y)/2$, linearization around the active-state fixed point $(\bar{\Sigma},\bar{\Delta}\neq 0)$ yields a non-normal Jacobian matrix:
$J=\begin{pmatrix}-\lambda_{1}&\omega_{\text{ff}}\\ 0&-\lambda_{2}\end{pmatrix},$ (2)
where the eigenvalues are given by $\lambda_{1}=\alpha+f(\bar{s})+(1-\bar{\Sigma})\omega_{0}f^{\prime}(\bar{s})$ and $\lambda_{2}=\alpha+f(\bar{s})$, with $\bar{s}=\omega_{0}\bar{\Sigma}+h$. The feedforward term $\omega_{\text{ff}}=(1-\bar{\Sigma})(\gamma_{\mu}+\gamma_{\nu})f^{\prime}(\bar{s})$ governs the asymmetric influence of $\Sigma$ on $\Delta$, while the parameter $\omega_{0}=\gamma_{\mu}-\gamma_{\nu}$ quantifies the difference between excitatory and inhibitory couplings. A key feature of this structure is that the eigenvectors of $J$ are non-orthogonal and tend to become nearly parallel under the balance condition $\zeta\equiv\frac{\gamma_{\mu}-\gamma_{\nu}}{\gamma_{\mu}+\gamma_{\nu}}\approx 0$. In this regime, the eigenbasis poorly spans the full phase space: directions orthogonal to the dominant eigenvectors are underrepresented, and any perturbation with a component in these directions requires large contributions from the available (nearly aligned) eigenvectors. As a result, even small imbalances along $\Delta$ can lead to disproportionately large transient excursions in $\Sigma$, despite the presence of negative eigenvalues and overall linear stability. This reflects the hallmark of non-normal dynamics: strong directional sensitivity and transient amplification that escape standard eigenvalue-based stability analysis.
This type of structure has important consequences, especially in the presence of stochasticity. In fact, Benayoun et al. introduced stochastic effects into the Wilson–Cowan framework by deriving it from a microscopic model with a finite population of $N$ neurons [23, 24, 13]. This approach leads to a system of coupled Langevin equations—interpreted in the Ito sense—that captures the impact of intrinsic fluctuations arising from finite-size effects:
$\frac{dx(t)}{dt}$ $=-\alpha x+(1-xf(s)+\sqrt{\alpha x+(1-x)f(s)}\eta_{x}(t)$ (3)
$\frac{dy(t)}{dt}$ $=-\alpha y+(1-y)f(s)+\sqrt{\alpha y+(1-y)f(s)}\eta_{y}(t)$
where $\eta_{x,y}$ are uncorrelated Gaussian white noises, with amplitude $\sigma$ that depends on the network-size $\sigma\propto 1/\sqrt{N}$ and some time dependencies have been omitted for simplicity in the notation.
A key insight is that even when the deterministic dynamics converge to a non-trivial stable fixed point (i.e., an active state), the presence of non-orthogonal, nearly degenerate eigenvectors can undermine this stability in the presence of noise [25, 13, 7] . Specifically, the poor representation of orthogonal directions in the eigen-basis makes the system highly sensitive to certain perturbations, allowing stochastic fluctuations to push it away from the active state and trap it near the quiescent one. As a consequence, the dynamics becomes intermittent, exhibiting avalanche-like activity: transient, noise-driven excursions of variable amplitude into higher activity states. This illustrates how non-normality, when combined with intrinsic stochasticity, can give rise to rich dynamical regimes that are entirely absent in the purely deterministic system (see [13, 7] and Section III.2 for further discussion).
Structural non-normal effects: feedforward and feedback network motifs.
To investigate how non-normal dynamics generalize from individual excitatory-inhibitory (x/y) units to interconnected circuits, we now analyze simple network motifs composed of multiple coupled x/y populations. Our goal is to understand how the non-normal effects identified at the single-unit level are shaped by the architecture of inter-unit interactions.
We focus on a minimal yet informative motif comprising three local x/y populations, as illustrated in Fig.1. Inter-population connectivity is mediated exclusively by excitatory neurons: these not only project to excitatory neurons in other units but also to inhibitory neurons across populations. In contrast, inhibitory neurons act only locally within each population, consistent with biological observations [26]. Because inter-population coupling is purely excitatory, we summarize the global architecture using the $\times 3$ 3 adjacency matrix $A$, which captures the excitatory-to-excitatory connectivity pattern:
$A=\begin{pmatrix}1&\epsilon&0\\ 0&1&\epsilon\\ c&0&1\end{pmatrix}$ (4)
The parameter $0\leq c\leq\epsilon$ interpolates between a symmetric cycle ($c=\epsilon$) and a purely feedforward structure ($c=0$), thus controlling the degree of non-normality. The full $\times 6$ 6 adjacency matrix, which includes both excitatory and inhibitory neurons, is not shown explicitly. However, it consists of this $\times 3$ 3 excitatory submatrix duplicated in the excitatory-to-inhibitory blocks, while the inhibitory subnetwork remains purely diagonal, since inhibition is strictly local.
Figure 1: Inter-populations interactions. a) Schematic representation of the generic non-reciprocal motif. X and Y respectively stand for the fraction of active excitatory and inhibitory populations. Excitatory couplings are represented in dark red, while inhibitory interactions are shown in dark green. Interactions between populations are illustrated by light green arrows. Two of these inter-population connections are modulated by the parameter $\epsilon$, while only one is modulated by the parameter $0\leq c\leq\epsilon$, according to Eq. 4. This allows us to control the behavior of the system between the homogeneous loop and the feed-forward case. b) Schematic representation of the model following from the transformation ($\Sigma$ = x+y, $\Delta$ = x-y). The inter-populations interactions are drawn in orange for the cyclic setup and in red for the feedforward one.
The full set of coupled Langevin equations that we consider is then:
$\begin{split}\dot{\text{x}}_{i}&=-\alpha\text{x}_{i}+(1-\text{x}_{i})f\bigg{(}h+A_{i,i}(\gamma_{\mu}\text{x}_{i}-\gamma_{\nu}\text{y}_{i})\\ &\hskip 136.5733pt+\gamma_{\ell}\sum_{j\neq i}A_{i,j}\text{x}_{j}\bigg{)}\\ \dot{\text{y}}_{i}&=-\alpha\text{y}_{i}+(1-\text{y}_{i})f\bigg{(}h+A_{i,i}(\gamma_{\mu}\text{x}_{i}-\gamma_{\nu}\text{y}_{i})\\ &\hskip 136.5733pt+\gamma_{\ell}\sum_{j\neq i}A_{i,j}\text{x}_{j}\bigg{)},\end{split}$ (5)
where $\gamma_{\ell}$ sets the excitatory coupling strength between units. To fix the ideas we will consider $f(s)=\Theta(s)\tanh{s}$, where $\Theta(s)$ is the Heaviside function, or its derivable counterpart $f_{\xi}(s)=\frac{1}{2}\left(1+\frac{2}{\pi}\arctan\left(\frac{s}{\xi}\right)\right)\tanh{s}$, with for $\xi<<1$).
We focus primarily on motifs composed of three x/y populations, which already capture rich and nontrivial dynamical behavior. Generalization to $n$ coupled populations is presented in the Supplemental Material (SM).
The Jacobian of this set of equations develops complex eigenvalues when three or more coupled units are considered, as detailed in Sec.III.2, allowing us one to account for more interesting transient and time-dependent behaviors such as noise-induced oscillations [27]. The paper is organized as follows: in Section III.1 we analyze the steady state of Eq. (5) for the cyclic and feedforward motifs; in Section III.2 we perform linear analysis to gain insight on the features of the orbits around the stable fixed point; in Section III.3 we examine the reactivity properties and in Section 4 we relate them to the behavior of the stochastic system.
III Results
III.1 Steady state analysis
First, we analyze the type of bifurcation that the deterministic system undergoes near the zero fixed point, i.e. when it shifts from inhibition-dominated to excitation-dominated, focusing on the comparison between the two limit architectures introduced above. Then, we study the controllability of the system, by quantifying how rapidly the network escapes from the inactive phase using the excitatory coupling $\gamma_{\mu}$ as a control parameter.
Figure 2: Bifurcation diagram and scaling of the fixed points. Panel a) shows the bifurcation diagram for basal population $x_{1}$ obtained via Runge-Kutta integration. The parameters $\alpha=0.1$, $\gamma_{\nu}=0.35$ and $\gamma_{\ell}=0.2$ have been used for several setups. The orange line confirms the bifurcation point at $\gamma_{\mu}=\gamma_{\mu}^{c,\epsilon}=0.25$ for the cyclic system. The decoupled diagram, with bifurcation point predicted at $\gamma_{\mu}=\gamma_{\mu}^{c}=0.45$, is shifted to make $\gamma_{\mu}^{c,\epsilon}=\gamma_{\mu}^{c}$. Remarkably, the cyclic and the decoupled lines coincide over the whole curve. Eventually, several values of $c$ are evaluated, up to $c=0$ (feed-forward setup), with expected bifurcation point $\gamma_{\mu}^{c,0}=0.45$. In the other panels, the scaling of the fixed point around the bifurcation. The behavior for $c=\epsilon/2$ in panels c), d) and e) shows how the linear scaling extends beyond $c=\epsilon$ with reduced but reliable approximation for each population of the cyclic setup. Panel b) refers to the feed-forward case.
For the sake of simplicity, let us start by discussing the decoupled setup ($\gamma_{\ell}=0$), studied in [13, 7]. In this case, the units are trivially just copies of the same system, i.e. $x_{i}=x_{j},y_{i}=y_{j}$ for all couples of indices $i,j$. More interestingly, the equilibrium points satisfy the identities: $\bar{x}_{i}=\bar{y}_{i}$, as anticipated above. We now would like to extend the validity of these equalities for a certain region around the fixed points. As we will discuss in detail in Section III.2, the system is strongly non-normal and both eigenvectors of the Jacobian align along the bisector of the $x_{i}-y_{i}$ plane. This allows us to suppose that for $t>>\infty$, $x_{i}(t)\sim y_{i}(t)$. Therefore, the dynamical system reduces to a one-dimensional single equation.
[…]
$\dot{\text{x}}=(-\alpha+\gamma_{\mu}-\gamma_{\nu})\text{x}-(\gamma_{\mu}-\gamma_{\nu})\text{x}^{2}+\mathcal{O}(x^{3}),$ (6)
which is the normal form for a transcritical bifurcation [28]. Then, the non-zero solution of this equation is
$\bar{\text{x}}\sim\frac{\gamma_{\mu}-\gamma_{\nu}-\alpha}{\gamma_{\mu}-\gamma_{\nu}}.$ (7)
and the bifurcation occurs at $\gamma_{\mu}^{c}=\gamma_{\nu}+\alpha$. Then, linearizing around this bifurcation point we get the scaling $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)$.
Cyclic coupling
In the cyclic architecture ($c=\epsilon$), the system’s symmetry ensures that the equations for each node are exactly the same. Then, the equilibrium condition $\bar{x}_{i}=\bar{x}_{j}$ is valid for all couple of indices $i,j$. Moreover, with a reasoning analogous to the one in [13], it is easy to show that also in this case the excitatory and inhibitory activities reach the same steady state $\bar{x}_{i}=\bar{y}_{i}$. Then, as in the decoupled case, the system allows a one-dimensional description.
The generalization of the decoupled system to the cyclic one turns out to be straightforward: it is sufficient to substitute $\gamma_{\mu}$ with $\gamma_{\mu}+\gamma_{\ell}$. Therefore, it follows that $\gamma_{\mu}^{c,\epsilon}=\gamma_{\nu}+\alpha-\gamma_{\ell}$, and the scaling of the non-null solution becomes $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c,\epsilon}\right)$. Fig. 2(a) and (b) report the bifurcation diagram for the population $x_{1}$ and the scaling of the fixed point around the bifurcation, showing that the curve for the cyclic systems as a function of $\gamma_{\mu}$ coincides with the curve for the decoupled system, once adequately shifted. Moreover, in Fig. 2 we show the scaling of $x_{1}$, as a function of the distance to the bifurcation point. Our numerical analyses reveal that the scaling exponent seems to hold not only for the cyclic and decoupled system, but also for a broad range of values of $c<\epsilon$, i.e. for architectures that are not perfect cycles, that we will refer to as ‘weak-feedback systems’.
This suggests that the cyclic interactions do not fundamentally alter the physics of the system, provided that feedback into $x_{3}$ is preserved. Thus, our bifurcation analysis suggests that the cyclic and decoupled models obey the same physics.
Feed-forward coupling
In the feedforward system, the units are organized in a chain. To fix the ideas, we label the different populations as follows: $x_{3}$ is the initial node of the chain and does not receive any input from other units, $x_{2}$ is in the middle of the chain receiving input from $x_{3}$ and projecting onto $x_{1}$, and $x_{1}$ does not project to any other units, being the final node of the chain (see Fig. 1 a)). The population $x_{3}$ trivially behaves as a decoupled unit, thus becoming active at the bifurcation point $\gamma_{\mu}=\gamma_{\mu}^{c}$. For smaller values of $\gamma_{\mu}$, $x_{2}$ receives no input from $x_{3}$ and therefore also $x_{2}$ stays inactive. The same is true also for $x_{1}$, therefore the whole system is inactive for $\gamma_{\mu}<\gamma_{\mu}^{c}$, while it gets triggered by the activation of $x_{3}$, whose input cascades onto the rest of the chain. More formally, around the bifurcation point, we find:
$\dot{\text{x}_{i}}=$ $(1-\bar{\text{x}}_{i+1})\bar{\mathrm{f}}-\alpha\text{x}_{i}-\bar{\mathrm{f}}\text{x}_{i}+\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}+$ (8)
$-\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}^{2}-\frac{1}{2}\bar{\mathrm{f}}^{\prime\prime}(\gamma_{\mu}-\gamma_{\nu})^{2}\text{x}_{i}^{2}+\mathcal{O}(x^{3}),$
where now $\mathrm{f}$ and its derivatives are no longer calculated at zero but in $\epsilon\gamma_{\ell}\bar{\text{x}}_{i+1}$, where $\bar{\text{x}}_{i+1}$ is an external parameter not affected by $\bar{\text{x}}_{i}$. Let us manipulate Eq. (8), to write it in the form $\dot{\text{x}_{i}}=c+b\text{x}_{i}+a\text{x}_{i}^{2}+\mathcal{O}(x^{3})$, by defining $\delta\equiv\gamma_{\mu}-\gamma_{\mu}^{c}$, $b=\delta$, $c=\frac{\gamma_{\nu}}{\alpha}\delta=\mathcal{O}(1)$ and $a=\frac{\gamma_{\nu}}{\alpha}\delta(\delta+\alpha)^{2}-(\delta+\alpha)=\mathcal{O}(1)$. It follows that for $\gamma_{\mu}>\gamma_{\nu}+\alpha$, Eq. (8) is formally equal to
$\dot{\text{x}_{i}}=(1-\bar{\text{x}}_{i+1})\bar{\mathrm{f}}-\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}^{2}-\frac{1}{2}\bar{\mathrm{f}}^{\prime\prime}(\gamma_{\mu}-\gamma_{\nu})^{2}\text{x}_{i}^{2}+\mathcal{O}(x^{3}),$ (9)
with solutions $\bar{\text{x}}\sim\pm\sqrt{\frac{c}{a}}$. This is the normal form for a saddle-node bifurcation [28].
Furthermore, using $\bar{\text{x}}_{3}=\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)+\mathcal{O}(\gamma_{\mu}^{2})$, we can verify that the bifurcation for $x_{2}$ occurs exactly at $\gamma_{\mu}=\gamma_{\mu}^{c}$, given $\bar{\mathrm{f}}\sim\epsilon\gamma_{\ell}\bar{\text{x}}_{3}$ and moreover
$\bar{\text{x}}_{2}\sim\frac{\sqrt{\epsilon\gamma_{\ell}}}{\alpha}\sqrt{\gamma_{\mu}-\gamma_{\mu}^{c}}.$ (10)
Then, by using Eq. (10), we can obtain the results for $x_{1}$: again we can verify that the bifurcation occurs at $\gamma_{\mu}^{c}$, and the scaling reads
$\bar{\text{x}}_{1}\sim\frac{(\epsilon\gamma_{\ell})^{3/4}}{\alpha}\sqrt[4]{\gamma_{\mu}-\gamma_{\mu}^{c}}$ (11)
Figure 2 (panels (a) and (c)) depicts the bifurcation diagram and the fixed points scaling, in agreement with our analytic findings. When we compare the scaling in the feedforward and cyclic cases, we can conclude that the fixed point of the feed-forward setup is more sensitive to $\gamma_{\mu}$ variations compared to the cyclic setup.
III.2 Non-normal dynamics in non-reciprocal motifs
In this section, first we focus on the calculation of eigenvectors of the Jacobian near the equilibrium points; next, we characterize the non-normality by defining a parameter that quantifies the angle between eigenvectors. This step will be key towards understanding the phenomenology of the stochastic regime. Again, we will focus on highlighting the differences between the cyclic and feedforward architectures.
Non-normal matrices (i.e. matrices that do not commute with their adjoint) typically have non-orthogonal eigenvectors and the angle between such eigenvectors is related to the non-normality of the system: the smaller the angle between two eigenvectors, the more prominent the effects of the non-normality.
Before proceeding with the linear analysis, it is useful to rewrite the system in terms of the following variables:
$\Sigma_{i}=\text{x}_{i}+\text{y}_{i}$ (12)
$\Delta_{i}=\text{x}_{i}-\text{y}_{i}.$
Since the solution will always be of the form $\text{x}_{i}=\text{y}_{i}$ $\forall i$, the equilibrium coordinate for the second variable will be $\bar{\Delta}_{i}=0$. Then, breaking down $\Sigma_{i}$ and $\Delta_{i}$ into the equilibrium and perturbation components, $\Sigma_{i}=\bar{\Sigma}_{i}+\sigma_{i}$ and $\Delta_{i}=\bar{\Delta}_{i}+\delta_{i}$, we can rewrite the linear expansion of the equations as follows
$\dot{\sigma}_{i}=-\alpha\sigma_{i}-\sigma\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma}_{i})\bar{\mathrm{f}}^{\prime}\biggl{[}\sigma(\gamma_{\mu}-\gamma_{\nu})+\delta(\gamma_{\mu}+\gamma_{\nu})+$ (13)
$\hskip 193.47873pt+\epsilon\gamma_{\ell}\frac{\sigma+\delta}{2}\biggr{]}$
$\dot{\delta}_{i}=-\alpha\delta_{i}-\delta\bar{\mathrm{f}}$
We observe that, in the new coordinate system, the Jacobian associated with the linear dynamics takes an upper triangular form. This change of basis, originally proposed in [13] for symmetry considerations in the decoupled system, effectively implements the Schur transformation of the Jacobian of 5, ensuring its upper triangular structure in the new variables.
The Jacobian matrices are calculated for generic fixed points since we are interested in the derivation of non-normality measures in generic regimes.
Cyclic couplings   Let’s begin with the study of the cyclic system ($c=\epsilon$). The Jacobian of the cyclic model has a block circulant matrix shape and thus it can be block diagonalized according to the circulant matrix theorem [29], as shown in SM. It is interesting to note that the block diagonalization leads to the basis $(\Sigma_{i},\Delta_{i})$. The same result can be achieved by a Schur transformation in the complex field.
A generic circulant matrix is diagonalized by a special unitary matrix known as the discrete Fourier transform matrix, which has $(r,s)$ element equal to $\exp(-2\pi\mathrm{i}(r-1)(s-1)/n)$. In our model, one can write the diagonalized Jacobian for the cyclic $J_{c}^{(D)}=SJ_{c}S^{-1}$ as follows:
$J_{c}^{(D)}=\begin{pmatrix}\Phi+\epsilon\Psi&0&0\\ 0&\Phi-\epsilon\Psi\biggl{(}\frac{1}{2}-i\frac{\sqrt{3}}{2}\biggr{)}&0\\ 0&0&\Phi-\epsilon\Psi\biggl{(}\frac{1}{2}+i\frac{\sqrt{3}}{2}\biggr{)}\\ \end{pmatrix}$ (14)
where
$\Phi=\begin{pmatrix}-\alpha-\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\gamma_{\nu})\\ 0&-\alpha-\bar{\mathrm{f}}\end{pmatrix},\hskip 8.53581pt$ (15)
$\psi=\begin{pmatrix}\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}\\ 0&0\end{pmatrix}$
Therefore, according to this base, the graph can be remapped into three disjoint feedforward graphs as depicted in Fig. 1b. Exploiting the upper triangular shape of the sub-matrices, we can read the eigenvalues in the diagonal elements:
$\lambda_{i}^{(1)}=$ $-\alpha-\bar{\mathrm{f}}$ (16)
$\lambda_{i}^{(2)}=$ $-\alpha-\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+k_{i}\gamma_{\ell}-\gamma_{\nu})$
where $k_{i}=\biggl{\{}\epsilon,\epsilon\biggl{(}-\frac{1}{2}+i\frac{\sqrt{3}}{2}\biggr{)},\epsilon\biggl{(}-\frac{1}{2}-i\frac{\sqrt{3}}{2}\biggr{)}\biggr{\}}$, $i=1,2,3$. We identify two complex and four real eigenvalues, wherein global stability is constrained by the inequality $-\alpha-\bar{f}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu})<0$. In particular the null fixed point loses its stability when $\gamma_{\mu}=\gamma_{\nu}+\alpha$, as already proved in Sec. III.1.
We are now ready to calculate the form of the eigenvectors. The full derivation, including a generalization to $n$ populations, can be found in SM. From the form of the eigenvectors’ elements, it becomes evident that $\xi=|{\gamma_{\nu}-\gamma_{\mu}-\epsilon\gamma_{\ell}}|/({\gamma_{\mu}+\gamma_{\nu}+\epsilon\gamma_{\ell}})$ controls the angle between two of the six eigenvectors. These two eigenvectors encompass contributions of excitatory and inhihbitory neurons across all populations; in particular one of them is defined by $\Sigma_{i}=1,\Delta_{i}=0$ –i.e. lies in the direction where all excitatory and inhibitory populations have the same firing rate– and the other one is defined by $\Sigma_{i}=1,\Delta_{i}=\xi$.
Thus, by tuning the parameter $\xi$ one can directly modulate the angle between these two eigenvectors, which determines the degree of the non-normality in the system.
In the asymptotic limit $\xi\rightarrow\infty$ the eigenvectors become orthogonal and the system becomes normal. Conversely, when $\xi=0$, $J_{c}^{(D)}$ becomes defective, as the two eigenvectors collapse into one. In this limit, non-normality emerges from a single structural mechanism, controlled exclusively by the parameter $\xi$. In other words, $\xi$ is the sole parameter through which the eigenvector collapse occurs. This parameter $\xi$ can be seen as a generalization of the parameter $\zeta={\gamma_{\nu}-\gamma_{\mu}}/({\gamma_{\mu}+\gamma_{\nu}})$ controlling non-normality in the decoupled system [7]. Furthermore, the degenerate eigenvector defined by $\Sigma_{i}=1,\Delta_{i}=0$ extends the corresponding structure found in the decoupled system.
Once again, by reabsorbing $\epsilon\gamma_{\ell}$ into $\gamma_{\mu}$ the cyclic system maps onto the decoupled one.
Feed-forward couplings   Let’s move on to the feedforward case by setting $c=0$. Now the Jacobian matrix takes the form:
$J_{f}=\begin{pmatrix}\Phi_{1}&\Psi_{1}\epsilon&0\\ 0&\Phi_{2}&\Psi_{2}\epsilon\\ 0&0&\Phi_{3}\\ \end{pmatrix},$ (17)
where the subscripts indicate that the 222x222 matrices $\Phi_{k}$ (respectively $\Psi_{k}$) share the same structure, but are evaluated at different equilibrium points, since each unit of the chain reaches a different steady state. However, as stated above, the equilibrium point of the excitatory and inhibitory sub-populations within each unit, remains homogeneous (i.e. $\bar{\text{x}}_{i}=\bar{\text{y}}_{i}$). The upper diagonal form of $J_{f}$ allows to read directly the eigenvalues on the diagonal (see SM), i.e.:
$\{-\alpha-\bar{\mathrm{f}}_{i},\frac{1}{2}(2-\bar{\Sigma}_{i})\bar{\mathrm{f}}^{\prime}_{i}(\gamma_{\mu}-\gamma_{\nu})\},\hskip 5.69054pti=1,2,3.$ (18)
Again, the complete form of the eigenvectors, along with their derivation and a generalization to $n$ populations, are provided in SM.
Interestingly, both the elements of the eigenvectors and the eigenvectors themselves exhibit a structured pattern that reflects a recursive-like organization, explicitly highlighting the directionality of signal propagation along the chain. Moreover, in this case, the non-normality parameter has two contributions. The first, encoding the so-called inner non-normality, is given by $\zeta=|{\gamma_{\nu}-\gamma_{\mu}}|/({\gamma_{\mu}+\gamma_{\nu}})$ which coincides with the non-normality parameter of the decoupled system [7]. The second contribution, representing the outer non-normality, is associated with the directionality of the flow along the chain, and coincides with the coupling parameter $\gamma_{\ell}$.
Similar to the cyclic setup, $J_{f}$ becomes defective for $\zeta\rightarrow 0$ and normal for $\zeta=\infty$. Additionally, in the feedforward system, the non-normality can also be influenced by $\gamma_{\ell}$. In this case, the Jacobian becomes defective also when $\gamma_{\ell}\rightarrow\infty$, while $\gamma_{\ell}$ is not able to alter the inner non-normality.
The parameters that we derived here to tune the angle between the eigenvectors, do not coincide with the Henrici measure, commonly used to quantify the strength of the non-normality [12]. A full comparison is discussed in SM.
Figure 3: Reactivity and comparison of numerical abscissa and non-normality parameters. On the right $\rho(t)$ (panels c), d), e) and f)) trajectories under reactivity. Reactivity increases from left to right with the numerical abscissa $m$. Here $\zeta$ represents the non-normality parameter for the binary system. Top row panels c) and d): In each panel the couplings are chosen for the decoupled system and remain unchanged over the setups. From c) and d) $\xi$ keeps fixed. Bottom row panels e) and f): As the setup changes, the non-normality of both the feedforward and cyclic systems is adjusted to maintain a consistent level across all configurations in each panel. With this change, only the feedforward system exhibits a significant increase in reactivity. On the left (panels a) and b)), the level curves of $m$ for different values of $\gamma_{\ell}$ are evaluated at the inactive phase. The colored areas correspond to the subspace where the level curves are meaningful. The straight lines correspond to the portions of the plane with fixed non-normality ($\zeta=cont$ or $\xi=cont$). The cyclic and the feedforward setups are respectively addressed in panels a) and b). In both cases, it can be observed that increasing $\gamma_{\ell}$ means shifting the $m$ level curves to the left. Then, for each set ($\gamma_{\nu}$,$\gamma_{\mu}$) $m$ is enhanced. As a final remark, there it appears clear how all the possible values of $m$ can be obtained while still fixing the non-normality parameter.
III.3 Reactivity
Reactivity is one of the distinctive markers of non-normality since it only arises in its presence. However, the reverse is not always true, i.e. not all non-normal systems exhibit reactivity. A reactive system is one in which, during a transient phase the system moves away from the steady state, following a perturbation. More specifically, at small time scales the 222-norm of a variable is enhanced before exponentially decaying back to the equilibrium point. Formally, reactivity occurs only when the maximum eigenvalue of the Hermitian part of the Jacobian, known as numerical abscissa $m$, is positive:
$m(\mathbf{J})=\sup\sigma\left[\left(\mathbf{J}+\mathbf{J}^{*}\right)/2\right],$ (19)
where $\sigma(\mathbf{J})$ denotes the spectrum of the matrix $\mathbf{J}$ and $\mathbf{J}^{*}$ is its conjugate transpose. The growth during the transient is proportional to the magnitude of $m(\mathbf{J})$ (as long as $m(\mathbf{J})>0$).
In the cyclic network, the numerical abscissa can be evaluated analytically. In particular, we find that for the relevant portion of parameter space $\gamma_{\mu},\gamma_{\nu}>0$, we obtain:
$m=$ $\frac{1}{2}(-2\alpha-2\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}))+$ (20)
$\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\sqrt{(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu})^{2}+(\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu})^{2}}$
The results are summarized in Fig. 3, where we present the level curves of $m$ for several values of $\gamma_{\ell}$ and for $c=0,\epsilon$. Notably, panel f) indicates that the behavior of the feed-forward system as a function of $m$ –whose analytic calculation is much more involved– qualitatively matches with that of the cyclic case, shown in panel e). This observation suggests that Eq.20 may serve as a qualitative reference even for intermediate configurations with $0<c<\epsilon$.
Interestingly, it turns out that reactivity is governed by two distinct quantities: $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}$ and $\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, with the absolute value of their ratio being the non-normality parameter. This implies that, although non-normality is needed to trigger reactivity, a fixed non-normality level does not correspond to a single level of reactivity. In other words, the ratio of $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}$ and $\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$ can remain constant while adjusting both numerator and denominator to enhance $m$. This is shown explicitly in Fig. 3 panels e) and f), where the lines representing constant values of $\xi$ or $\zeta$ cross the level curves of $m$, i.e., while moving along a line with constant non-normality, multiple values of reactivity can be achieved.
Moreover, from Eq. 20 we define $\Gamma\equiv\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, i.e. the sum of all coupling constants. In what follows we fix the non-normality and distinguish between two regimes: one with weak interactions and weak reactivity ($\Gamma,m\ll 1$, illustrated in Fig. 3 panels c), e)) and one of strong interactions and strong reactivity ($\Gamma,m\gg 1$, illustrated in panels d), f)).
In the right dashed region, we show the time-evolution of the total activity $\rho(t)$ $=\sqrt{\sum_{i}x_{i}^{2}+y^{2}_{i}}$, starting from an initial condition close to the bisector, where Eq. 20 holds. In the first row, parameters are chosen so that the inner systems are all equivalent, and in particular, the non-normality parameter of the inner systems, $\zeta$, is kept constant. We want to compare the evolution of the norm $\rho(t)$ when the reactivity is small (panel c)) with a case in which reactivity is large (panel d)). As expected we observe that, when the reactivity is large, the system moves away from the fixed point in 0 before relaxing. From these plots, the cyclic system seems to be more reactive than the decoupled system, and that reactivity increases further as $c$ approaches $\epsilon$.
However, the reactivity may appear larger for the coupled systems, simply because the coupling itself increases the non-normality, which has an effect in the reactivity. To isolate the effects of reactivity, in the lower row, we fix the non-normality of each system to the same value. In particular we choose parameters in the cyclic system such that $\xi=\zeta$, i.e. we adjust $\gamma_{\mu}$ to compensate for the introduction of $\gamma_{\ell}$.
Again, the left panel, panel e) is for small values of reactivity, and the right panel, panel f) is for larger values of the reactivity, showing larger excursions away from the fixed point. Two main messages emerge from these plots: first, we verify that the cyclic system can be reduced to the decoupled case merely through coupling parameter shifts; second, we show that the feedforward system undergoes a genuine increase in reactivity when internal non-normality is held constant.
For the feedforward system, instead, we do not have a unique global parameter to quantify the total non-normality and summing the parameters for inner and outer non-normality would be an arbitrary choice. A specific choice of the couplings that makes the reactivity coincide surely exists, but estimating it is not easy. The point that we want to make here is that the external source of non-normality enhances the reactivity.
In conclusion, the feed-forward network actually increases reactivity compared to the decoupled system. In contrast, in the cyclic system the enhancement of reactivity can be fully accounted for by equalizing the level of non-normality.
III.4 Analysis of the stochastic evolution of the dynamics
As shown in [13, 7], the microscopic underlying system shows very strong finite-size effects. In particular it was shown in [7] using Eq. 3, that the demographic noise induces a minimum of the effective potential is the origin, thus generating a bistability in the system. In presence of strong non-normality, the shear flow along the diagonal, associated with the two nearly-collapsing eigenvectors, facilitates the jumps between the two fixed points and the dynamics resembles the avalanching behaviors observed in neural system, where the global activity undergoes large irregular fluctuations between a low-activity and a high-activity state.
Here we want to explore the behavior of the coupled system in response to noise. We then couple three mean-field dynamics, where each obeys the system size expansion as in [13]:
$\frac{dx_{i}}{dt}$ $=-\alpha x_{i}+(1-x_{i})f(s_{i})+\sqrt{\alpha x_{i}+(1-x_{i})f(s_{i})}\eta_{x_{i}}$ (21)
$\frac{dy_{i}}{dt}$ $=-\alpha y_{i}+(1-y_{i})f(s_{i})+\sqrt{\alpha y_{i}+(1-y_{i})f(s_{i})}\eta_{y_{i}}.$
Also in this case, the introduction of noise induces a new stable equilibrium point in zero for all parameter choices, given that $h$ is kept small. Again, when the deterministic stable point bifurcates from zero, bistability appears. Bistability is not a sufficient ingredient to observe avalanching behavior: a high non-normality is needed [7]. However, for too high values of $\xi$ and $\zeta$ the system tends to be trapped in the active minimum. Since we are interested in how the activity propagates across populations, we used the fraction of active excitatory neurons to visualize the trajectories, Fig. 4. Thanks to the non-normality, similar plots could be obtained for inhibitory activities.
For low values of the inner non-normality, the system easily gets stuck in its minimum, and excursions between the two minima are not likely.
Once a sufficiently large non-normality level is set, we can distinguish between two separate behaviours, one where the trajectories are incoherent and one where the system shows dynamical patterns, as a function of the parameter $\gamma_{\ell}$. When $\gamma_{\ell}<<\gamma_{\mu},\gamma_{\nu}$ the system lies in an incoherent phase (Fig. 4 panels b) and d)), while when $\gamma_{\ell}\sim\gamma_{\mu},\gamma_{\nu}$ the behavior is more coherent (Fig. 4 panels a) and c)). When the connectivity among units is large enough and the architecture is cyclic, the activation between units is highly correlated (Fig. 4a)). Instead, in the feed-forward case, when the connectivity among units is large enough, the units are activated sequentially. To fix the ideas, let us take two adjacent populations, where the bottom one projects input to –but does not receive any input from– the top one. When the bottom population transits to the active minimum, it activates also the upper population. The process propagates along the chain, giving rise to the step-like trajectories we observe in Fig. 4c). A discussion about the effective potential induced by the noise and the external couplings is detailed in SM.
Figure 4: Comparison of cyclic and feedforward setups in the stochastic regime. The top panels illustrate the cyclic configuration, while the bottom ones refer to the feedforward setup. The comparison between the left and the right columns shows how the system can be driven from an incoherent regime (on the right, panels b) and d)) to a coherent phase in THE cyclic setup (panel a)) and to step-like trajectories in the feed-forward one (panel c)). Parameter choices are as follows: a) $\gamma_{\ell}=1.13$ and b) $\gamma_{\ell}=0.13$; for both $\xi=0.004$. c) $\gamma_{\ell}=1.13$ and d) $\gamma_{\ell}=0.07$; for both $\zeta=0.005$.
IV Discussion and conclusions
Firstly, through the analysis of bifurcations and reactivity, we find that systems with feedback are more controllable, meaning that changes in parameter settings – or changes in physiological conditions – result in smoother alterations of network responses compared to purely feedforward systems. This enhanced controllability implies that the activation patterns of neurons –which are the substrate for neural representation of the stimulus– are more robust to perturbations when feedback is present. Similarly, it has been proposed that when stimuli become consciously reportable, they elicit an ‘ignition process’, which causes information about a brief stimulus to become sustained and broadcasted across many brain areas [30]. Our work suggests, in an extremely simplified setup, that the presence of feedback might be a key ingredient in such process. More specifically, our model reveals that the introduction of feedback into the system plays a critical role in generating coherence between processing stages. This observation aligns with experimental findings of ‘zero-lag interaction’ between layers during perception and visual working memory [31, 32]. These are frequently attributed to either common external inputs or specific network motifs [31, 33]. Interestingly, our model suggests that feedback alone might suffice to produce this synchrony, even in the absence of external modulation. Exploring the influence of time-lagged feedback would be a valuable future direction to test the validity of this hypothesis.
Additionally, our model shows that in a purely feedforward setup, the stochastic dynamics exhibits a greater degree of variability and complexity in the system’s response patterns, compared to the cyclic setup. In the absence of feedback, the neural activity is more distributed across a broader range of activation patterns, reflecting a more diverse set of responses to sensory inputs.
Remarkably, when the inter-unit connectivity is weak, coherent representations are completely lost, both for feedforward and cyclic networks.
A priori, we chose $n=3$ to investigate the potential role of complex eigenvalues and to explore whether the system could sustain quasi-cycles under the influence of noise. However, our analysis reveals that, despite the theoretical presence of complex eigenvalues, no oscillatory behavior emerges in practice. As $n$ increases, the frequency of the oscillation decreases, and the quasi-cycle dynamics potentially predicted by the mathematical formalism might be observable for larger $n$. This could be investigated with numerical simulations of systems with $n>3$, but is beyond the scope of the present work.
This study contributes to the growing body of literature suggesting that feedback plays a pivotal role in the neural dynamics underlying perception. Our model provides new insights into how feedback mechanisms could inherently promote synchrony and robustness in neural networks, even in the absence of common input.

### v2

@@
 Abstract
-Non-reciprocal interactions are a defining feature of many complex systems, biological, ecological, and technological, often pushing them far from equilibrium and enabling rich dynamical responses.
-These asymmetries can arise at multiple levels: locally, in the dynamics of individual units, and globally, in the topology of their interactions.
-In this work, we investigate how these two forms of non-reciprocity interact in networks of neuronal populations.
-At the local level, each population is modeled by a non-reciprocally coupled set of excitatory and inhibitory neural populations exhibiting transient amplification and reactivity.
-At the network level, these populations are coupled via directed, asymmetric connections that introduce structural non-normality.
-Since non-reciprocal interactions generically lead to non-normal linear operators, we frame both local and global asymmetries in terms of non-normal dynamics.
-Using a modified Wilson–Cowan framework, we analyze how the interplay between these two types of non-normality shapes the system’s behavior.
-We show that their combination leads to emergent collective dynamics, including fluctuation-driven transitions, dimensionality reduction, and novel nonequilibrium steady states.
-Our results provide a minimal yet flexible framework to understand how multi-scale non-reciprocities govern complex dynamics in neural and other interconnected systems.
+Non-reciprocal interactions are widespread in neural circuits, arising both from asymmetric local excitation–inhibition dynamics and from directed coupling between distinct populations.
+Here we investigate how these two sources of asymmetry jointly shape network activity using a modified Wilson-Cowan framework in which each node is an excitatory/inhibitory population operating in a strongly non-normal regime and nodes are linked by directed excitatory projections.
+We analyze minimal three-node motifs that interpolate between a feedback (cyclic) architecture and a purely feedforward chain.
+By combining steady-state and bifurcation analysis with linear stability and reactivity measures, we show that feedback coupling largely maps onto an effective parameter shift relative to the uncoupled case, whereas feedforward coupling produces qualitatively different activation properties, including distinct scaling across stages and a genuine enhancement of reactive transient amplification.
+Incorporating intrinsic finite-size noise, we find fluctuation-driven switching between inactive and active states and motif-dependent spatiotemporal organization: feedback promotes coherent, correlated activation across nodes, while feedforward networks display sequential, step-like propagation.
+Together, these results provide a minimal framework for understanding how multi-scale non-reciprocities control transient responses and noise-driven dynamics in coupled neural populations.
 I Introduction
@@
 Non-reciprocal interactions break detailed balance and time-reversal symmetry at the microscopic level, enabling new types of collective dynamics —including, among others, spontaneous oscillations, traveling waves, and fluctuation-amplified responses— that do not exist in symmetric, equilibrium-like systems [31, 18, 3, 22].
-In this paper, we aim to bring together and clarify two distinct, but often coexisting, manifestations of non-reciprocity in complex systems, both of which can be naturally interpreted through the lens of non-normality: a mathematical property of matrices where eigenvectors are non-orthogonal, allowing for transient amplification of perturbations.
-While non-normality provides a technical language to describe these effects, we emphasize that it is the underlying physical non-reciprocity — whether in the local dynamics or the network architecture — that drives the phenomena of interest.
-The first form we address is dynamical non-normality, which arises at the level of the internal dynamics of individual units or subsystems.
-When the local Jacobian or stability matrix is non-normal, even stable fixed points can exhibit large transient responses to small perturbations, especially in the presence of noise [12, 13].
-In neural circuits, for instance, populations of excitatory and inhibitory neurons often operate near critical balance, leading to strongly non-normal dynamics that transiently amplify fluctuations in input, contributing to fast signal processing, selectivity, and flexible population coding [14, 15, 16].
-These effects are intrinsically linked to asymmetries in the internal coupling and can be viewed as a form of microscopic non-reciprocity between excitatory and inhibitory neurons.
-The second form is what we term structural non-normality, which stems from the network-level asymmetry in the way local populations are connected at a larger scale.
-In many real-world systems, especially biological and engineered networks, the connectivity matrix that governs inter-unit influences is inherently non-reciprocal.
-Neural circuits, for example, are not fully bidirectional: higher-order cortical areas send feedback to early sensory regions, but not necessarily with equal weight or latency [17, 18, 19].
-Similarly, ecological and social systems often exhibit trophic or influence hierarchies, where some magnitude flows in preferred directions [20].
-When such asymmetries are strong, the overall network becomes structurally non-normal, leading to directional instabilities, long-range correlations, waves, and flow-like dynamics.
-The central hypothesis we aim to explore is whether the interplay between local dynamical non-normality and global structural non-normality gives rise to qualitatively new collective behaviors, beyond what either form could produce in isolation.
-In other words, when locally non-reciprocal units (e.g., excitatory/inhibitory local neural populations) are embedded in a globally non-reciprocal network (e.g., feedforward/feedback cortical circuits), the system exhibits emergent properties that are not captured by standard linear response theory or symmetric interaction models.
-These might include enhanced stochastic amplification, novel transient modes, and selective propagation of fluctuations and signals across the network.
-As a test of this hypothesis, here we analyze a minimal yet general neural model consisting of multiple interacting dynamical units, each described by a non-normal local system, coupled through a network, that in general is non-reciprocal.
-Using tools from linear stability analysis, stochastic dynamics, and spectral theory, we explore how the composition of non-reciprocities across scales governs the system’s susceptibility to noise, its capacity to transmit signals, and its repertoire of emergent dynamical states.
+In this paper, we distinguish two manifestations of physical non-reciprocity in complex systems—arising either from asymmetric local excitatory-inhibitory dynamics at a local level or from directed network architecture—which often co-occur but are conceptually separate.
+We use non-normal linear operators as a convenient mathematical framework to describe both types of non-reciprocity and analyze their dynamical consequences.
+Recall that a matrix is non-normal when it does not commute with its transpose (for real matrices) or adjoint (for complex matrices), e.g.
+$AA^{\top}\neq A^{\top}A$.
+In such cases, its (right) eigenmodes are generally non-orthogonal, and this non-orthogonality can enable strong transient amplification even when all eigenvalues indicate asymptotic stability.
+The first form we address is local dynamical non-normality, which arises at the level of the internal dynamics of individual units or subsystems.
+For example, consider an island with a predator population and a prey population.
+The local Jacobian (stability matrix) around a stationary point of the dynamics can be non-normal, so that even if the fixed point is asymptotically stable, the system may exhibit large transient responses to small perturbations—an effect that can be enhanced by continual disturbances or noise [30, 29, 35, 42, 37, 36].
+Similarly, in neural circuits, interacting excitatory and inhibitory neuronal populations often operate in regimes where the effective linearized dynamics is strongly non-normal, allowing transient amplification of input fluctuations and supporting fast signal processing, selectivity, and flexible population coding [19, 2, 3, 1, 27, 12, 6, 39, 4].
+The second form is what we term structural non-normality, which stems from network-level asymmetry in how units influence one another at a larger scale.
+In many real-world systems, especially biological, and engineered networks, the connectivity that mediates inter-unit interactions is directed and typically not symmetric.
+Ecological and social systems, for example, often exhibit trophic or influence hierarchies, in which some quantity flows preferentially along directed pathways [20, 23].
+Neural circuits likewise are not fully bidirectional: higher-order cortical areas send feedback to early sensory regions, but not necessarily with equal strength in the opposite direction [17, 14, 32, 18].
+Such directed structure can (but need not) generate a non-normal connectivity operator; when it does, it can promote directional amplification, long-range correlations, wave-like responses, and flow-like propagation of activity across the network.
+Here, we test the hypothesis that embedding locally non-normal units within globally non-reciprocal architectures yields collective behaviors stronger or distinct from those produced by either source of asymmetry alone.
+Specifically, we explore how the interplay between local dynamical non-normality (within E-I populations) and global structural non-normality (across network tiers) determines the reactivity of the system.
+We show that while feedback architectures largely rescale the effective parameters of local units, feedforward motifs fundamentally alter the scaling laws of the system, introducing a hierarchy of timescales and amplifying transient responses.
+In other words, when locally asymmetric (and potentially non-normal) subsystems (e.g., excitatory/inhibitory microcircuits) are embedded in a directed large-scale architecture (e.g., feedforward/feedback pathways), the resulting macroscopic dynamics may exhibit emergent phenomena not captured by standard linear-response intuition for symmetric interactions.
+These may include enhanced stochastic amplification, novel transient modes, and selective propagation of fluctuations and signals across the network.
+As a test of this hypothesis, here we analyze a minimal yet general neural model consisting of multiple interacting dynamical units, each described by a non-normal local E-I system, coupled through a network, that in general is non-reciprocal.
+Using tools from linear stability analysis, stochastic dynamics, and spectral theory, we explore how the composition of non-reciprocities —dynamical and structural— governs the system’s susceptibility to noise, its capacity to transmit signals, and its repertoire of emergent dynamical states.
 While our focus is primarily on neural systems —where both forms of non-reciprocity are particularly notorious— the conceptual framework we propose is broadly applicable to other domains, including active materials, biological networks, and collective decision-making systems.
-In all these contexts, non-reciprocity serves as a dynamical and structural engine for complexity, and understanding its layered manifestations is key to unlocking the behavior of high-dimensional, complex systems.
-A theoretical neuroscience perspective:
+In all these settings, non-reciprocity functions as a structural and dynamical source of complexity, and elucidating its distinct manifestations is crucial for understanding the behavior of high-dimensional, complex systems.
 II Wilson-Cowan model for neural populations
-Dynamical non-normal effects at a single region
+II.1 Dynamical non-normal effects in E/I populations.
 We consider the classical Wilson-Cowan model [40], which describes the dynamics of a large neural population composed of homogeneously interconnected excitatory and inhibitory neurons [10, 41, 5, 12].
 The system of differential equations governing the dynamical evolution of the model is given by:
-[…]
-where $x(t)$ and $y(t)$ are the activation levels of the excitatory and inhibitory neuronal populations, respectively.
-The parameter $\alpha$ represents the rate of spontaneous activity decay, while $s$ denotes the net incoming current, $s(t)=\gamma_{\mu}x(t)-\gamma_{\nu}y(t)+h$, which is simply the sum of all synaptic excitatory and inhibitory inputs, weighted by their respective synaptic strengths —-$(\gamma_{\mu}$ for excitation and $\gamma_{\nu}$ for inhibition— plus a constant external input current, $h$.
-To reduce the dimensionality of the parameter space, all synaptic weights of a given type have been set to the same value.
-Finally $f(s)$ is a sigmoid non-negative response function.
-The system displays two distinct dynamical regimes: a quiescent (inactive) phase, where $\bar{x}=\bar{y}=0$, and an active phase $\bar{x}=\bar{y}>0$, characterized by nonzero steady-state activity with excitation dominating the response ($\bar{x},\bar{y}\neq 0$).
+$\begin{aligned} \frac{dx(t)}{dt}&=-\alpha x(t)+(1-x(t))f(s)\\ \frac{dy(t)}{dt}&=-\alpha y(t)+(1-y(t))f(s),\end{aligned}$ (1)
+where $x(t)$ and $y(t)$ represent the levels of activation of excitatory and inhibitory neuronal population, respectively.
+The parameter $\alpha$ represents the rate of spontaneous activity decay, while $s$ denotes the net incoming input, $s(t)=\gamma_{\mu}x(t)-\gamma_{\nu}y(t)+h$, which is simply the sum of all synaptic excitatory and inhibitory inputs, weighted by their respective synaptic strengths plus a constant external input, $h$.
+To reduce the dimensionality of the parameter space, all synaptic weights of a given type have been set to the same value: $(\gamma_{\mu}$ for excitation and $\gamma_{\nu}$ for inhibition).
+Finally, $f(s)$ is a sigmoid non-negative response function.
+The autonomous system ($h=0$) displays two distinct dynamical regimes: a quiescent (inactive) phase, where the mean activity vanishes, $\bar{x}=\bar{y}=0$, and an active phase characterized by nonzero steady-state activity $\bar{x}=\bar{y}>0$.
 These phases are separated by a bifurcation that occurs at the critical point $\gamma_{\mu}=\gamma_{\nu}+\alpha$: below this threshold, the system remains in the quiescent state, while above it, activity becomes self-sustained (see below).
-In the deterministic limit, using variables $\Sigma=(x+y)/2$ and $\Delta=(x-y)/2$, linearization around the active-state fixed point $(\bar{\Sigma},\bar{\Delta}\neq 0)$ yields a non-normal Jacobian matrix:
+It is useful to make the change of variables $\Sigma=(x+y)/2$ and $\Delta=(x-y)/2$, to consider the overall activation and the differential one between excitation and inhibition, respectively.
+Linearizing around the active-state fixed point $(\bar{\Sigma},\bar{\Delta}\neq 0)$, one obtains a strictly upper-triangular Jacobian matrix, which is manifestly non-normal:
 $J=\begin{pmatrix}-\lambda_{1}&\omega_{\text{ff}}\\ 0&-\lambda_{2}\end{pmatrix}$, (2)
@@
 The feedforward term $\omega_{\text{ff}}=(1-\bar{\Sigma})(\gamma_{\mu}+\gamma_{\nu})f^{\prime}(\bar{s})$ governs the asymmetric influence of $\Sigma$ on $\Delta$, while the parameter $\omega_{0}=\gamma_{\mu}-\gamma_{\nu}$ quantifies the difference between excitatory and inhibitory couplings.
-A key feature of this structure is that the eigenvectors of $J$ are non-orthogonal and tend to become nearly parallel under the balance condition $\zeta\equiv\frac{\gamma_{\mu}-\gamma_{\nu}}{\gamma_{\mu}+\gamma_{\nu}}\approx 0$.
-In this regime, the eigenbasis poorly spans the full phase space: directions orthogonal to the dominant eigenvectors are underrepresented, and any perturbation with a component in these directions requires large contributions from the available (nearly aligned) eigenvectors.
-As a result, even small imbalances along $\Delta$ can lead to disproportionately large transient excursions in $\Sigma$, despite the presence of negative eigenvalues and overall linear stability.
+A key feature of this structure is that the eigenvectors of $J$ are non-orthogonal [5, 12] and tend to become nearly parallel under the condition $\zeta\equiv\frac{|\gamma_{\mu}-\gamma_{\nu}|}{\gamma_{\mu}+\gamma_{\nu}}\approx 0$, which describes balance between excitation and inhibition.
+In this regime, the eigenbasis poorly spans the full phase space: directions orthogonal to the dominant eigenvectors are poorly represented, and any perturbation with a component in these directions requires large contributions from the available (nearly aligned) eigenvectors.
+As a result, even small fluctuations along $\Delta$ can lead to giant transient excursions in $\Sigma$, despite the presence of negative eigenvalues and overall linear stability [5, 12, 8].
 This reflects the hallmark of non-normal dynamics: strong directional sensitivity and transient amplification that escape standard eigenvalue-based stability analysis.
@@
 This approach leads to a system of coupled Langevin equations —interpreted in the Ito sense— that captures the impact of intrinsic fluctuations arising from finite-size effects:
-$\frac{dx(t)}{dt}$ $=-\alpha x+(1-xf(s)+\sqrt{\alpha x+(1-x)f(s)}\eta_{x}(t)$ (3)
+$\frac{dx(t)}{dt}$ $=-\alpha x+(1-x)f(s)+\sqrt{\alpha x+(1-x)f(s)}\eta_{x}(t)$ (3)
 $\frac{dy(t)}{dt}$ $=-\alpha y+(1-y)f(s)+\sqrt{\alpha y+(1-y)f(s)}\eta_{y}(t)$
-where $\eta_{x,y}$ are uncorrelated Gaussian white noises, with amplitude $\sigma$ that depends on the network-size $\sigma\propto 1/\sqrt{N}$ and some time dependencies have been omitted for simplicity in the notation.
+where $\eta_{x,y}$ are uncorrelated Gaussian white noises, with amplitude $\sigma$ that depends on the network-size $\sigma\propto 1/\sqrt{N}$ (note that some time dependencies have been omitted for simplicity in the notation).
+Figure 1: Inter-population interactions.
+(a): Schematic representation of the generic non-reciprocal motif.
+$X$ and $Y$ denote the fractions of active excitatory and inhibitory populations, respectively.
+Local excitatory couplings are shown in dark green, while inhibitory interactions are indicated in dark red.
+Light green arrows illustrate inter-population interactions.
+Two of these connections (represented in light blue) are modulated by the parameter $c$, with $0\leq c\leq\epsilon$, as defined in Eq.
+(4).
+This modulation allows to interpolate between the homogeneous loop and the feedforward configuration.
+(b): Schematic representation of the model in the transformed variables ($\Sigma=x+y$, $\Delta=x-y$).
+In the transformed variables (b) the feedforward configuration includes both within-unit (orange) and between-units (red) couplings, while the cyclic network only includes within-unit (orange) interactions, i.e.
+the three units are disjoint.
 A key insight is that even when the deterministic dynamics converge to a non-trivial stable fixed point (i.e., an active state), the presence of non-orthogonal, nearly degenerate eigenvectors can undermine this stability in the presence of noise [28, 5, 12] .
@@
 This illustrates how non-normality, when combined with intrinsic stochasticity, can give rise to rich dynamical regimes that are entirely absent in the purely deterministic system (see [5, 12] and Section III.2 for further discussion).
-Structural non-normal effects: feedforward and feedback network motifs.
-To investigate how non-normal dynamics generalize from individual excitatory-inhibitory (x/y) units to interconnected circuits, we now analyze simple network motifs composed of multiple coupled x/y populations.
-Our goal is to understand how the non-normal effects identified at the single-unit level are shaped by the architecture of inter-unit interactions.
-We focus on a minimal yet informative motif comprising three local x/y populations, as illustrated in Fig.1.
-Inter-population connectivity is mediated exclusively by excitatory neurons: these not only project to excitatory neurons in other units but also to inhibitory neurons across populations.
-In contrast, inhibitory neurons act only locally within each population, consistent with biological observations [26].
-Because inter-population coupling is purely excitatory, we summarize the global architecture using the $\times 3$ 3 adjacency matrix $A$, which captures the excitatory-to-excitatory connectivity pattern:
+II.2 Structural non-normal effects: feedforward and feedback network motifs.
+To investigate how non-normal dynamics extend from individual excitatory-inhibitory (x/y) units to interconnected circuits, we analyze simple network motifs composed of multiple coupled x/y populations.
+Our aim is to understand how the non-normal features identified at the single-unit level are modulated by the architecture of inter-unit interactions.
+Here we focus on motifs composed of three x/y populations (or "units" or "areas"), schematically illustrated in Fig.
+1, which already capture rich and nontrivial dynamical behavior (generalization to $n$ coupled populations and more complex motifs are analyzed in the Supplemental Material).
+Inter-population connectivity is mediated exclusively by excitatory neurons projecting to both excitatory and inhibitory neurons in other populations.
+In contrast, inhibitory neurons act only locally within each population, in accordance with biological observations [21].
+Because inter-population coupling is purely excitatory, the global architecture can be compactly described by a $3\times 3$ adjacency matrix $A$, which encodes the excitatory-to-excitatory connectivity pattern:
 $A=\begin{pmatrix}1&\epsilon&0\\ 0&1&\epsilon\\ c&0&1\end{pmatrix}$ (4)
-The parameter $0\leq c\leq\epsilon$ interpolates between a symmetric cycle ($c=\epsilon$) and a purely feedforward structure ($c=0$), thus controlling the degree of non-normality.
-The full $\times 6$ 6 adjacency matrix, which includes both excitatory and inhibitory neurons, is not shown explicitly.
-However, it consists of this $\times 3$ 3 excitatory submatrix duplicated in the excitatory-to-inhibitory blocks, while the inhibitory subnetwork remains purely diagonal, since inhibition is strictly local.
-Figure 1: Inter-populations interactions.
-a) Schematic representation of the generic non-reciprocal motif.
-X and Y respectively stand for the fraction of active excitatory and inhibitory populations.
-Excitatory couplings are represented in dark red, while inhibitory interactions are shown in dark green.
-Interactions between populations are illustrated by light green arrows.
-Two of these inter-population connections are modulated by the parameter $\epsilon$, while only one is modulated by the parameter $0\leq c\leq\epsilon$, according to Eq.
-4.
-This allows us to control the behavior of the system between the homogeneous loop and the feed-forward case.
-b) Schematic representation of the model following from the transformation ($\Sigma$ = x+y, $\Delta$ = x-y).
-The inter-populations interactions are drawn in orange for the cyclic setup and in red for the feedforward one.
-The full set of coupled Langevin equations that we consider is then:
+The parameter $0\leq c\leq\epsilon$ interpolates between a symmetric cycle ($c=\epsilon$) and a purely feedforward structure ($c=0$), thereby controlling the degree of non-normality.
+The full $6\times 6$ adjacency matrix, which includes both excitatory and inhibitory neurons, is not shown explicitly (but note that it is constructed from the previous $3\times 3$ excitatory submatrix: this submatrix is duplicated in the excitatory-to-inhibitory blocks, while the inhibitory subnetwork remains purely diagonal, reflecting the strictly local nature of inhibition).
+The full set of coupled Wilson-Cowan equations that we consider is then:
 $\begin{split}\dot{\text{x}}_{i}&=-\alpha\text{x}_{i}+(1-\text{x}_{i})f\bigg(h+A_{i,i}(\gamma_{\mu}\text{x}_{i}-\gamma_{\nu}\text{y}_{i})\\ &\hskip 136.5733pt+\gamma_{\ell}\sum_{j\neq i}A_{i,j}\text{x}_{j}\bigg)\\ \dot{\text{y}}_{i}&=-\alpha\text{y}_{i}+(1-\text{y}_{i})f\bigg(h+A_{i,i}(\gamma_{\mu}\text{x}_{i}-\gamma_{\nu}\text{y}_{i})\\ &\hskip 136.5733pt+\gamma_{\ell}\sum_{j\neq i}A_{i,j}\text{x}_{j}\bigg),\end{split}$ (5)
-where $\gamma_{\ell}$ sets the excitatory coupling strength between units.
-To fix the ideas we will consider $f(s)=\Theta(s)\tanh{s}$, where $\Theta(s)$ is the Heaviside function, or its derivable counterpart $f_{\xi}(s)=\frac{1}{2}\left(1+\frac{2}{\pi}\arctan\left(\frac{s}{\xi}\right)\right)\tanh{s}$, with for $\xi<<1$).
-We focus primarily on motifs composed of three x/y populations, which already capture rich and nontrivial dynamical behavior.
-Generalization to $n$ coupled populations is presented in the Supplemental Material (SM).
-The Jacobian of this set of equations develops complex eigenvalues when three or more coupled units are considered, as detailed in Sec.III.2, allowing us one to account for more interesting transient and time-dependent behaviors such as noise-induced oscillations [27].
-The paper is organized as follows: in Section III.1 we analyze the steady state of Eq.
-(5) for the cyclic and feedforward motifs; in Section III.2 we perform linear analysis to gain insight on the features of the orbits around the stable fixed point; in Section III.3 we examine the reactivity properties and in Section 4 we relate them to the behavior of the stochastic system.
+with $i=,1,2$ or $3$ and where $\gamma_{\ell}$ sets the excitatory coupling strength between units, and $f(s)=\Theta(s)\tanh{s}$, with $\Theta(s)$ the Heaviside function.
+The Jacobian of this system develops complex eigenvalues when three or more coupled units are considered, as detailed in Sec.
+III.2, allowing us to account for interesting transient and time-dependent behaviors, such as noise-induced oscillations [13].
+These $\epsilon$-dependent minimal motifs allow us to isolate how architectural directionality modifies bifurcation structure, eigenvector geometry, reactivity, and noise-driven dynamics, while keeping the local non-normal mechanism fixed.
+The rest of the paper is organized as follows: in Sec.
+III.1 we analyze the steady state of Eq.
+(5) for the cyclic and feedforward motifs; in Sec.
+III.2 we perform a linear stability analysis to characterize the features of the orbits around the stable fixed point; in Sec.
+III.3 we examine the reactivity properties; and in Sec.
+III.4 we relate these properties to the behavior of the stochastic system.
 III Results
 III.1 Steady state analysis
-First, we analyze the type of bifurcation that the deterministic system undergoes near the zero fixed point, i.e.
-when it shifts from inhibition-dominated to excitation-dominated, focusing on the comparison between the two limit architectures introduced above.
-Then, we study the controllability of the system, by quantifying how rapidly the network escapes from the inactive phase using the excitatory coupling $\gamma_{\mu}$ as a control parameter.
+First, we analyze the type of bifurcation that the deterministic system undergoes near the vanishing-activity fixed point, i.e., when the system transitions from inhibition-dominated to excitation-dominated dynamics, focusing on the comparison between the two limit architectures introduced above.
+Subsequently, we study the controllability of the system by quantifying how abruptly the network escapes from the inactive phase, using the excitatory coupling $\gamma_{\mu}$ as a control parameter.
 Figure 2: Bifurcation diagram and scaling of the fixed points.
-Panel a) shows the bifurcation diagram for basal population $x_{1}$ obtained via Runge-Kutta integration.
-The parameters $\alpha=0.1$, $\gamma_{\nu}=0.35$ and $\gamma_{\ell}=0.2$ have been used for several setups.
-The orange line confirms the bifurcation point at $\gamma_{\mu}=\gamma_{\mu}^{c,\epsilon}=0.25$ for the cyclic system.
-The decoupled diagram, with bifurcation point predicted at $\gamma_{\mu}=\gamma_{\mu}^{c}=0.45$, is shifted to make $\gamma_{\mu}^{c,\epsilon}=\gamma_{\mu}^{c}$.
-Remarkably, the cyclic and the decoupled lines coincide over the whole curve.
-Eventually, several values of $c$ are evaluated, up to $c=0$ (feed-forward setup), with expected bifurcation point $\gamma_{\mu}^{c,0}=0.45$.
-In the other panels, the scaling of the fixed point around the bifurcation.
-The behavior for $c=\epsilon/2$ in panels c), d) and e) shows how the linear scaling extends beyond $c=\epsilon$ with reduced but reliable approximation for each population of the cyclic setup.
-Panel b) refers to the feed-forward case.
-For the sake of simplicity, let us start by discussing the decoupled setup ($\gamma_{\ell}=0$), studied in [13, 7].
-In this case, the units are trivially just copies of the same system, i.e.
-$x_{i}=x_{j},y_{i}=y_{j}$ for all couples of indices $i,j$.
-More interestingly, the equilibrium points satisfy the identities: $\bar{x}_{i}=\bar{y}_{i}$, as anticipated above.
-We now would like to extend the validity of these equalities for a certain region around the fixed points.
-As we will discuss in detail in Section III.2, the system is strongly non-normal and both eigenvectors of the Jacobian align along the bisector of the $x_{i}-y_{i}$ plane.
-This allows us to suppose that for $t>>\infty$, $x_{i}(t)\sim y_{i}(t)$.
-Therefore, the dynamical system reduces to a one-dimensional single equation.
-[…]
+Panels (a–c) Scaling of the fixed points around the bifurcation for $x_{1},x_{2}$, and $x_{3}$ respectively.
+The scaling is the same for the decoupled case (purple), for $c=\epsilon/2$ (light blue) and $c=\epsilon$ (orange).
+(d) Scaling of the fixed points around the bifurcation for the Feedforward case ($c=0$).
+Note the change of the slope as expected by theoretical results, depending on the population (see legend).
+(e) Bifurcation diagram for the basal population $x_{1}$, obtained via Runge-Kutta integration.
+Parameters $\alpha=0.1$, $\gamma_{\nu}=0.35$, and $\gamma_{\ell}=0.2$ are used across several setups.
+The orange line marks the bifurcation point at $\gamma_{\mu}=\gamma_{\mu}^{c,\epsilon}=0.25$ for the cyclic system.
+The decoupled diagram, with bifurcation point predicted at $\gamma_{\mu}=\gamma_{\mu}^{c}=0.45$, is shifted so that $\gamma_{\mu}^{c,\epsilon}=\gamma_{\mu}^{c}$, resulting in a near-perfect overlap of the cyclic and decoupled curves.
+Several values of $c$ are also evaluated, up to $c=0$ (feedforward setup), with the expected bifurcation point $\gamma_{\mu}^{c,0}=0.45$.
+For the sake of simplicity, we begin by discussing the decoupled setup ($\gamma_{\ell}=0$), previously studied in [5, 12].
+In this case, the units are trivially copies of the same system, i.e., $x_{i}=x_{j}$ and $y_{i}=y_{j}$ for all pairs of indices $i,j$ and the equilibrium points satisfy $\bar{x}_{i}=\bar{y}_{i}$.
+We now aim to extend the validity of these equalities to a region surrounding the fixed points.
+As discussed in detail in Sec.
+III.2, the system is strongly non-normal, and both eigenvectors of the Jacobian align along the bisector of the $x_{i}$$y_{i}$ plane.
+This alignment allows us to assume that, for $t\gg 1$, $x_{i}(t)\sim y_{i}(t)$.
+Consequently, the dynamical system effectively reduces to a one-dimensional equation.
+In particular, the normal form around the bifurcation point $\text{x}=0$, with $h=0$, can be written as
 $\dot{\text{x}}=(-\alpha+\gamma_{\mu}-\gamma_{\nu})\text{x}-(\gamma_{\mu}-\gamma_{\nu})\text{x}^{2}+\mathcal{O}(x^{3}),$ (6)
-which is the normal form for a transcritical bifurcation [28]. Then, the non-zero solution of this equation is
+which formally corresponds to the normal form for a transcritical bifurcation [34]. The non-zero solution of this equation is
 $\bar{\text{x}}\sim\frac{\gamma_{\mu}-\gamma_{\nu}-\alpha}{\gamma_{\mu}-\gamma_{\nu}}.$ (7)
-and the bifurcation occurs at $\gamma_{\mu}^{c}=\gamma_{\nu}+\alpha$.
-Then, linearizing around this bifurcation point we get the scaling $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)$.
-Cyclic coupling
-In the cyclic architecture ($c=\epsilon$), the system’s symmetry ensures that the equations for each node are exactly the same.
-Then, the equilibrium condition $\bar{x}_{i}=\bar{x}_{j}$ is valid for all couple of indices $i,j$.
-Moreover, with a reasoning analogous to the one in [13], it is easy to show that also in this case the excitatory and inhibitory activities reach the same steady state $\bar{x}_{i}=\bar{y}_{i}$.
-Then, as in the decoupled case, the system allows a one-dimensional description.
-The generalization of the decoupled system to the cyclic one turns out to be straightforward: it is sufficient to substitute $\gamma_{\mu}$ with $\gamma_{\mu}+\gamma_{\ell}$.
-Therefore, it follows that $\gamma_{\mu}^{c,\epsilon}=\gamma_{\nu}+\alpha-\gamma_{\ell}$, and the scaling of the non-null solution becomes $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c,\epsilon}\right)$.
+and the bifurcation occurs at the critical value $\gamma_{\mu}^{c}=\gamma_{\nu}+\alpha$.
+Then, linearizing around this bifurcation point one gets $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)$.
+III.1.1 Cyclic coupling
+In the cyclic architecture ($c=\epsilon$), the system’s symmetry ensures that the equations for each node are identical.
+Consequently, the equilibrium condition $\bar{x}_{i}=\bar{x}_{j}$ holds for all pairs of indices $i,j$.
+Moreover, following a reasoning as the one above, one can show that, as before, excitatory and inhibitory activities also reach the same steady state, $\bar{x}_{i}=\bar{y}_{i}$.
+Therefore, as in the decoupled case, the dynamics can be effectively described by a one-dimensional equation.
+The generalization of the decoupled system to the cyclic architecture is straightforward: it is sufficient to replace $\gamma_{\mu}$ with $\gamma_{\mu}+\gamma_{\ell}$.
+Consequently, the bifurcation point becomes
+$\gamma_{\mu}^{c,\epsilon}=\gamma_{\nu}+\alpha-\gamma_{\ell},$ (8)
+and the scaling of the non-zero solution is
+$\bar{x}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c,\epsilon}\right),$ (9)
+i.e., as in the decoupled case, the level of activity grows linearly with the distance to the bifurcation point (thus, the corresponding critical exponent is $\beta=1).$
 Fig.
-2(a) and (b) report the bifurcation diagram for the population $x_{1}$ and the scaling of the fixed point around the bifurcation, showing that the curve for the cyclic systems as a function of $\gamma_{\mu}$ coincides with the curve for the decoupled system, once adequately shifted.
-Moreover, in Fig.
-2 we show the scaling of $x_{1}$, as a function of the distance to the bifurcation point.
-Our numerical analyses reveal that the scaling exponent seems to hold not only for the cyclic and decoupled system, but also for a broad range of values of $c<\epsilon$, i.e.
-for architectures that are not perfect cycles, that we will refer to as ‘weak-feedback systems’.
-This suggests that the cyclic interactions do not fundamentally alter the physics of the system, provided that feedback into $x_{3}$ is preserved.
-Thus, our bifurcation analysis suggests that the cyclic and decoupled models obey the same physics.
-Feed-forward coupling
+2(a) and (b) show the bifurcation diagram for the population $x_{1}$ and the scaling of the fixed point around the bifurcation point, demonstrating that the curve for the cyclic system as a function of $\gamma_{\mu}$ coincides with that of the decoupled system once appropriately shifted (see figure caption).
+Moreover, the figure illustrates the scaling of $x_{1}$ as a function of the distance to the bifurcation point, with an exponent $\beta=1$.
+The figure also illustrates that this scaling exponent holds not only for the cyclic and decoupled systems, but also for a broad range of values $c<\epsilon$, corresponding to architectures that are not perfect cycles, which we refer to as "weak-feedback systems."
+These results suggest that cyclic interactions do not fundamentally alter the phase transition of the system, with respect to the uncoupled case, provided that feedback is preserved ($c>0$).
+III.1.2 Purely feedforward coupling
 In the feedforward system, the units are organized in a chain.
-To fix the ideas, we label the different populations as follows: $x_{3}$ is the initial node of the chain and does not receive any input from other units, $x_{2}$ is in the middle of the chain receiving input from $x_{3}$ and projecting onto $x_{1}$, and $x_{1}$ does not project to any other units, being the final node of the chain (see Fig.
-1 a)).
-The population $x_{3}$ trivially behaves as a decoupled unit, thus becoming active at the bifurcation point $\gamma_{\mu}=\gamma_{\mu}^{c}$.
-For smaller values of $\gamma_{\mu}$, $x_{2}$ receives no input from $x_{3}$ and therefore also $x_{2}$ stays inactive.
-The same is true also for $x_{1}$, therefore the whole system is inactive for $\gamma_{\mu}<\gamma_{\mu}^{c}$, while it gets triggered by the activation of $x_{3}$, whose input cascades onto the rest of the chain.
-More formally, around the bifurcation point, we find:
-$\dot{\text{x}_{i}}=$ $(1-\bar{\text{x}}_{i+1})\bar{\mathrm{f}}-\alpha\text{x}_{i}-\bar{\mathrm{f}}\text{x}_{i}+\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}+$ (8)
+We label the populations such that $x_{3}$ is the upstream or initial node, projecting to $x_{2}$, which in turn projects to $x_{1}$.
+The population $x_{3}$ behaves as a decoupled unit, becoming active at the bifurcation point $\gamma_{\mu}=\gamma_{\mu}^{c}$, and for $\gamma_{\mu}\geq\gamma_{\mu}^{c}$, activation of $x_{3}$ triggers a cascade.
+Around the bifurcation point, the dynamics of unit $i$ driven by unit $i+1$ can be expanded as:
+$\dot{\text{x}_{i}}$ $\approx\bar{\mathrm{f}}-\alpha\text{x}_{i}-\bar{\mathrm{f}}\text{x}_{i}+\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}$ (10)
 $-\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}^{2}-\frac{1}{2}\bar{\mathrm{f}}^{\prime\prime}(\gamma_{\mu}-\gamma_{\nu})^{2}\text{x}_{i}^{2}+\mathcal{O}(x^{3}),$
-where now $\mathrm{f}$ and its derivatives are no longer calculated at zero but in $\epsilon\gamma_{\ell}\bar{\text{x}}_{i+1}$, where $\bar{\text{x}}_{i+1}$ is an external parameter not affected by $\bar{\text{x}}_{i}$.
-Let us manipulate Eq.
-(8), to write it in the form $\dot{\text{x}_{i}}=c+b\text{x}_{i}+a\text{x}_{i}^{2}+\mathcal{O}(x^{3})$, by defining $\delta\equiv\gamma_{\mu}-\gamma_{\mu}^{c}$, $b=\delta$, $c=\frac{\gamma_{\nu}}{\alpha}\delta=\mathcal{O}(1)$ and $a=\frac{\gamma_{\nu}}{\alpha}\delta(\delta+\alpha)^{2}-(\delta+\alpha)=\mathcal{O}(1)$.
-It follows that for $\gamma_{\mu}>\gamma_{\nu}+\alpha$, Eq.
-(8) is formally equal to
-$\dot{\text{x}_{i}}=(1-\bar{\text{x}}_{i+1})\bar{\mathrm{f}}-\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}^{2}-\frac{1}{2}\bar{\mathrm{f}}^{\prime\prime}(\gamma_{\mu}-\gamma_{\nu})^{2}\text{x}_{i}^{2}+\mathcal{O}(x^{3}),$ (9)
-with solutions $\bar{\text{x}}\sim\pm\sqrt{\frac{c}{a}}$. This is the normal form for a saddle-node bifurcation [28].
-Furthermore, using $\bar{\text{x}}_{3}=\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)+\mathcal{O}(\gamma_{\mu}^{2})$, we can verify that the bifurcation for $x_{2}$ occurs exactly at $\gamma_{\mu}=\gamma_{\mu}^{c}$, given $\bar{\mathrm{f}}\sim\epsilon\gamma_{\ell}\bar{\text{x}}_{3}$ and moreover
-$\bar{\text{x}}_{2}\sim\frac{\sqrt{\epsilon\gamma_{\ell}}}{\alpha}\sqrt{\gamma_{\mu}-\gamma_{\mu}^{c}}.$ (10)
-Then, by using Eq. (10), we can obtain the results for $x_{1}$: again we can verify that the bifurcation occurs at $\gamma_{\mu}^{c}$, and the scaling reads
-$\bar{\text{x}}_{1}\sim\frac{(\epsilon\gamma_{\ell})^{3/4}}{\alpha}\sqrt[4]{\gamma_{\mu}-\gamma_{\mu}^{c}}$ (11)
-Figure 2 (panels (a) and (c)) depicts the bifurcation diagram and the fixed points scaling, in agreement with our analytic findings.
-When we compare the scaling in the feedforward and cyclic cases, we can conclude that the fixed point of the feed-forward setup is more sensitive to $\gamma_{\mu}$ variations compared to the cyclic setup.
+where $\bar{\mathrm{f}}$ and its derivatives are evaluated for the averaged external input $s=\epsilon\gamma_{\ell}\bar{\text{x}}_{i+1}$ (except for $i=3$, that has no input).
+By defining $\delta\equiv\gamma_{\mu}-\gamma_{\mu}^{c}$, we map this to the normal form $\dot{x}=c+ax^{2}$.
+The constant driving term scales as $c\propto\bar{\mathrm{f}}\propto\bar{x}_{i+1}$.
+Since $x_{3}$ scales linearly with $\delta$ (standard transcritical scaling), $x_{2}$ undergoes a saddle-node bifurcation driven by $x_{3}$, yielding:
+$\bar{\text{x}}_{2}\sim\frac{\sqrt{\epsilon\gamma_{\ell}}}{\alpha}\sqrt{\gamma_{\mu}-\gamma_{\mu}^{c}}\sim\delta^{1/2}.$ (11)
+Recursively, $x_{1}$ is driven by $x_{2}$, leading to a fourth-root scaling:
+$\bar{\text{x}}_{1}\sim\frac{(\epsilon\gamma_{\ell})^{3/4}}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)^{1/4}\sim\delta^{1/4}.$ (12)
+Generalizing this recursive argument to a feedforward chain of length $N$, the fixed point of the $k$-th node in the chain (where $k=1$ is the input and $k=N$ is the output) must scale as:
+$\bar{x}_{k}\sim(\gamma_{\mu}-\gamma_{\mu}^{c})^{1/2^{k-1}},$ (13)
+i.e.
+the exponent $\beta$ depends one the depth of the activated area.
+This result implies a rapid "flattening" of the response curve as one moves deeper into the chain.
+For deep feedforward circuits, the downstream nodes become effectively binary —switching sharply between inactive and active states— thereby acting as robust signal rectifiers despite the linear nature of the initial upstream perturbation.
+Fig.
+2(a/b) show the bifurcation diagram and the scaling of the fixed points in numerical simulations, revealing a perfect agreement with analytical predictions.
+Comparing the feedforward and cyclic cases, we observe that the fixed point in the feedforward setup is more sensitive to variations in $\gamma_{\mu}$ than in the cyclic architecture.
 III.2 Non-normal dynamics in non-reciprocal motifs
-In this section, first we focus on the calculation of eigenvectors of the Jacobian near the equilibrium points; next, we characterize the non-normality by defining a parameter that quantifies the angle between eigenvectors.
-This step will be key towards understanding the phenomenology of the stochastic regime.
-Again, we will focus on highlighting the differences between the cyclic and feedforward architectures.
-Non-normal matrices (i.e.
-matrices that do not commute with their adjoint) typically have non-orthogonal eigenvectors and the angle between such eigenvectors is related to the non-normality of the system: the smaller the angle between two eigenvectors, the more prominent the effects of the non-normality.
-Before proceeding with the linear analysis, it is useful to rewrite the system in terms of the following variables:
+Non-normal matrices, i.e.
+matrices that do not commute with their adjoint (the transpose in the real case), typically have non-orthogonal eigenvectors.
+In particular, the angle between their right eigenvectors (measured with the standard Euclidean inner product) provides a convenient proxy for non-normality: the smaller the angle, the more pronounced the associated transient effects.
+Motivated by this, we analyze the eigenvectors of the Jacobian near the equilibrium points and introduce a parameter that quantifies the angle between right eigenvectors.
+This characterization is key to understanding the phenomenology of the stochastic regime and to highlighting the differences between the cyclic and feedforward architectures.
+Let us first rewrite the system in terms of the previously defined variables, $\Sigma$ and $\Delta$
 $\Sigma_{i}=\text{x}_{i}+\text{y}_{i}$ (14)
 $\Delta_{i}=\text{x}_{i}-\text{y}_{i}.$
-Since the solution will always be of the form $\text{x}_{i}=\text{y}_{i}$ $\forall i$, the equilibrium coordinate for the second variable will be $\bar{\Delta}_{i}=0$.
-Then, breaking down $\Sigma_{i}$ and $\Delta_{i}$ into the equilibrium and perturbation components, $\Sigma_{i}=\bar{\Sigma}_{i}+\sigma_{i}$ and $\Delta_{i}=\bar{\Delta}_{i}+\delta_{i}$, we can rewrite the linear expansion of the equations as follows
-$\dot{\sigma}_{i}=-\alpha\sigma_{i}-\sigma\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma}_{i})\bar{\mathrm{f}}^{\prime}\biggl{[}\sigma(\gamma_{\mu}-\gamma_{\nu})+\delta(\gamma_{\mu}+\gamma_{\nu})+$ (13)
-$\hskip 193.47873pt+\epsilon\gamma_{\ell}\frac{\sigma+\delta}{2}\biggr{]}$
-$\dot{\delta}_{i}=-\alpha\delta_{i}-\delta\bar{\mathrm{f}}$
-We observe that, in the new coordinate system, the Jacobian associated with the linear dynamics takes an upper triangular form.
-This change of basis, originally proposed in [13] for symmetry considerations in the decoupled system, effectively implements the Schur transformation of the Jacobian of 5, ensuring its upper triangular structure in the new variables.
-The Jacobian matrices are calculated for generic fixed points since we are interested in the derivation of non-normality measures in generic regimes.
-Cyclic couplings   Let’s begin with the study of the cyclic system ($c=\epsilon$).
-The Jacobian of the cyclic model has a block circulant matrix shape and thus it can be block diagonalized according to the circulant matrix theorem [29], as shown in SM.
-It is interesting to note that the block diagonalization leads to the basis $(\Sigma_{i},\Delta_{i})$.
+Since the solution is always of the form $\text{x}_{i}=\text{y}_{i}$ $\forall i$, then $\bar{\Delta}_{i}=0$.
+Therefore, breaking down $\Sigma_{i}$ and $\Delta_{i}$ into the equilibrium and perturbation components, $\Sigma_{i}=\bar{\Sigma}_{i}+\sigma_{i}$ and $\Delta_{i}=\bar{\Delta}_{i}+\delta_{i}$, one can rewrite the linear expansion of the equations as follows:
+$\dot{\sigma}_{i}=-(\alpha+\bar{\mathrm{f}})\sigma_{i}+(1-\bar{\Sigma}_{i})\bar{\mathrm{f}}^{\prime}\biggl[\sigma_{i}(\gamma_{\mu}-\gamma_{\nu})+\delta_{i}(\gamma_{\mu}+\gamma_{\nu})+$ (15)
+$\hskip 165.02597pt+\epsilon\gamma_{\ell}(\sigma_{i-1}+\delta_{i-1})\biggr]$
+$\dot{\delta}_{i}=-(\alpha+\bar{\mathrm{f}})\delta_{i}.$
+Note that, in the new coordinate system, the Jacobian associated with the linearized dynamics assumes an upper-triangular form.
+This change of basis, originally proposed in [5] for symmetry considerations in the decoupled system, effectively implements the Schur transformation of the Jacobian of Eq.
+(5), guaranteeing its upper-triangular structure in the transformed variables.
+The Jacobian matrices are computed for generic fixed points, as our goal is to derive measures of non-normality applicable across regimes.
+III.2.1 Cyclic couplings
+Let us begin with the study of the cyclic system ($c=\epsilon$).
+The Jacobian of the cyclic model has a block circulant matrix shape and thus can be block diagonalized according to the circulant matrix theorem [11], as shown in Supp.
+Mat.
+A.1.
+It is interesting to note that the diagonalization of the blocks leads to the basis $(\Sigma_{i},\Delta_{i})$.
 The same result can be achieved by a Schur transformation in the complex field.
 A generic circulant matrix is diagonalized by a special unitary matrix known as the discrete Fourier transform matrix, which has $(r,s)$ element equal to $\exp(-2\pi\mathrm{i}(r-1)(s-1)/n)$.
-In our model, one can write the diagonalized Jacobian for the cyclic $J_{c}^{(D)}=SJ_{c}S^{-1}$ as follows:
+In our model, one can write the diagonalized Jacobian for the cyclic $J_{c}^{(D)}$ as follows:
 $J_{c}^{(D)}=\begin{pmatrix}\Phi+\epsilon\Psi&0&0\\ 0&\Phi-\epsilon\Psi\biggl(\frac{1}{2}-i\frac{\sqrt{3}}{2}\biggr)&0\\ 0&0&\Phi-\epsilon\Psi\biggl(\frac{1}{2}+i\frac{\sqrt{3}}{2}\biggr)\\ \end{pmatrix}$ (16)
 where
-$\Phi=\begin{pmatrix}-\alpha-\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\gamma_{\nu})\\ 0&-\alpha-\bar{\mathrm{f}}\end{pmatrix},\hskip 8.53581pt$ (15)
-$\psi=\begin{pmatrix}\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}\\ 0&0\end{pmatrix}$
-Therefore, according to this base, the graph can be remapped into three disjoint feedforward graphs as depicted in Fig.
+[…]
+$\Psi=\begin{pmatrix}\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}.\\ 0&0\end{pmatrix}$.
+Therefore, according to this base, the neural circuit can be remapped into three disjoint feedforward graphs as depicted in Fig.
 1b.
-Exploiting the upper triangular shape of the sub-matrices, we can read the eigenvalues in the diagonal elements:
+Using the upper triangular shape of the sub-matrices, we can calculate the eigenvalues in the diagonal elements:
 $\lambda_{i}^{(1)}=$ $-\alpha-\bar{\mathrm{f}}$ (18)
@@
 where $k_{i}=\biggl\{\epsilon,\epsilon\biggl(-\frac{1}{2}+i\frac{\sqrt{3}}{2}\biggr),\epsilon\biggl(-\frac{1}{2}-i\frac{\sqrt{3}}{2}\biggr)\biggr\}$, $i=1,2,3$.
-We identify two complex and four real eigenvalues, wherein global stability is constrained by the inequality $-\alpha-\bar{f}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu})<0$.
-In particular the null fixed point loses its stability when $\gamma_{\mu}=\gamma_{\nu}+\alpha$, as already proved in Sec.
+We can identify two complex and four real eigenvalues, wherein global stability is constrained by the inequality $\alpha-\bar{f}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu})<0$.
+In particular the null fixed point loses its stability when $\gamma_{\mu}=\gamma_{\nu}+\alpha$, as already found in Sec.
 III.1.
-We are now ready to calculate the form of the eigenvectors.
-The full derivation, including a generalization to $n$ populations, can be found in SM.
-From the form of the eigenvectors’ elements, it becomes evident that $\xi=|{\gamma_{\nu}-\gamma_{\mu}-\epsilon\gamma_{\ell}}|/({\gamma_{\mu}+\gamma_{\nu}+\epsilon\gamma_{\ell}})$ controls the angle between two of the six eigenvectors.
-These two eigenvectors encompass contributions of excitatory and inhihbitory neurons across all populations; in particular one of them is defined by $\Sigma_{i}=1,\Delta_{i}=0$ –i.e.
+Computing the eigenvectors’ elements it becomes evident —-as explicitly shown in in Supp.
+Mat.
+A.1.— that $\xi=|{\gamma_{\nu}-\gamma_{\mu}-\epsilon\gamma_{\ell}}|/({\gamma_{\mu}+\gamma_{\nu}+\epsilon\gamma_{\ell}})$ controls the angle between two of the six right eigenvectors.
+These two eigenvectors encompass contributions of excitatory and inhibitory neurons across all populations; in particular one of them is defined by $\Sigma_{i}=1,\Delta_{i}=0$ –i.e.
 lies in the direction where all excitatory and inhibitory populations have the same firing rate– and the other one is defined by $\Sigma_{i}=1,\Delta_{i}=\xi$.
-Thus, by tuning the parameter $\xi$ one can directly modulate the angle between these two eigenvectors, which determines the degree of the non-normality in the system.
-In the asymptotic limit $\xi\rightarrow\infty$ the eigenvectors become orthogonal and the system becomes normal.
-Conversely, when $\xi=0$, $J_{c}^{(D)}$ becomes defective, as the two eigenvectors collapse into one.
-In this limit, non-normality emerges from a single structural mechanism, controlled exclusively by the parameter $\xi$.
-In other words, $\xi$ is the sole parameter through which the eigenvector collapse occurs.
-This parameter $\xi$ can be seen as a generalization of the parameter $\zeta={\gamma_{\nu}-\gamma_{\mu}}/({\gamma_{\mu}+\gamma_{\nu}})$ controlling non-normality in the decoupled system [7].
+Thus, by tuning the parameter $\xi$ one can directly modulate the angle between these two (right) eigenvectors, which determines the degree of non-normality in the system.
+In the asymptotic limit $\xi\to\infty$, the eigenvectors become orthogonal and the system approaches normality.
+Conversely, for $\xi=0$, $J_{c}^{(D)}$ becomes defective, as the two eigenvectors coalesce into one.
+In this regime, non-normality stems from a single structural mechanism governed solely by $\xi$, which therefore uniquely controls the eigenvector collapse.
+From this perspective, $\xi$ can be viewed as a generalization of the parameter $\zeta={|\gamma_{\nu}-\gamma_{\mu}|}/({\gamma_{\mu}+\gamma_{\nu}})$, which quantifies non-normality in the decoupled system [12].
 Furthermore, the degenerate eigenvector defined by $\Sigma_{i}=1,\Delta_{i}=0$ extends the corresponding structure found in the decoupled system.
 Once again, by reabsorbing $\epsilon\gamma_{\ell}$ into $\gamma_{\mu}$ the cyclic system maps onto the decoupled one.
-Feed-forward couplings   Let’s move on to the feedforward case by setting $c=0$. Now the Jacobian matrix takes the form:
+Importantly, this result shows that feedback alone does not generate an independent source of structural non-normality: as long as cyclic symmetry is preserved, the collective dynamics remain reducible to those of a single effective E/I unit.
+This provides a useful baseline against which genuinely non-reducible architectures, such as feedforward chains, can be identified.
+III.2.2 Feedforward couplings
+We now move on to the feedforward case by setting $c=0$. In this case, the Jacobian matrix takes the form:
 $J_{f}=\begin{pmatrix}\Phi_{1}&\Psi_{1}\epsilon&0\\ 0&\Phi_{2}&\Psi_{2}\epsilon\\ 0&0&\Phi_{3}\\ \end{pmatrix}$, (19)
-where the subscripts indicate that the 222x222 matrices $\Phi_{k}$ (respectively $\Psi_{k}$) share the same structure, but are evaluated at different equilibrium points, since each unit of the chain reaches a different steady state.
+where the subscripts indicate that the 22x22 matrices $\Phi_{k}$ (respectively $\Psi_{k}$) share the same structure, but are evaluated at different equilibrium points, since each unit of the chain reaches a different steady state.
 However, as stated above, the equilibrium point of the excitatory and inhibitory sub-populations within each unit, remains homogeneous (i.e.
 $\bar{\text{x}}_{i}=\bar{\text{y}}_{i}$).
-The upper diagonal form of $J_{f}$ allows to read directly the eigenvalues on the diagonal (see SM), i.e.:
+The upper diagonal form of $J_{f}$ allows to read directly the eigenvalues on the diagonal (see Supp.
+Mat.
+A.2), i.e.:
 $\{-\alpha-\bar{\mathrm{f}}_{i},\frac{1}{2}(2-\bar{\Sigma}_{i})\bar{\mathrm{f}}^{\prime}_{i}(\gamma_{\mu}-\gamma_{\nu})\},\hskip 5.69054pti=1,2,3.$ (20)
-Again, the complete form of the eigenvectors, along with their derivation and a generalization to $n$ populations, are provided in SM.
+Again, the complete form of the eigenvectors, along with their derivation and a generalization to $n$ populations, are provided in Supp. Mat. A.2.
 Interestingly, both the elements of the eigenvectors and the eigenvectors themselves exhibit a structured pattern that reflects a recursive-like organization, explicitly highlighting the directionality of signal propagation along the chain.
-Moreover, in this case, the non-normality parameter has two contributions.
+By comparing the pairs of eigenvectors in the Supp.
+Mat.
+A.2, two contributions to the system’s non-normality can be identified.
 The first, encoding the so-called inner non-normality, is given by $\zeta=|{\gamma_{\nu}-\gamma_{\mu}}|/({\gamma_{\mu}+\gamma_{\nu}})$ which coincides with the non-normality parameter of the decoupled system [12].
-The second contribution, representing the outer non-normality, is associated with the directionality of the flow along the chain, and coincides with the coupling parameter $\gamma_{\ell}$.
-Similar to the cyclic setup, $J_{f}$ becomes defective for $\zeta\rightarrow 0$ and normal for $\zeta=\infty$.
-Additionally, in the feedforward system, the non-normality can also be influenced by $\gamma_{\ell}$.
-In this case, the Jacobian becomes defective also when $\gamma_{\ell}\rightarrow\infty$, while $\gamma_{\ell}$ is not able to alter the inner non-normality.
-The parameters that we derived here to tune the angle between the eigenvectors, do not coincide with the Henrici measure, commonly used to quantify the strength of the non-normality [12].
-A full comparison is discussed in SM.
+Similar to the cyclic setup, $J_{f}$ becomes defective for $\zeta\rightarrow 0$ and normal in the limit $\zeta\rightarrow\infty$.
+Specifically, if $\zeta=0$ the eigenvectors in Supp.
+Mat.
+A.2., collapse in pairs.
+The second contribution, representing the outer non-normality, is associated with the directionality of the flow along the chain, and coincides with the coupling parameter
+$\gamma_{\ell}$
+.
+If, in addition to $\zeta=0$, the inter-layer coupling is strong $(\gamma_{\ell}\gg\gamma_{\mu},\gamma_{\nu})$, then all six eigenvectors nearly collapse, and the feedforward architecture yields a dramatic amplification of non-normality.
+However, if $\zeta\neq 0$, then no pairs of eigenvectors collapse: $J_{f}$ cannot become defective by the effect of the directed network architecture alone.
 Figure 3: Reactivity and comparison of numerical abscissa and non-normality parameters.
-On the right $\rho(t)$ (panels c), d), e) and f)) trajectories under reactivity.
+Left panels.
+(a,b): Level curves of the numerical abscissa $m$ for different values of $\gamma_{\ell}$, evaluated at the inactive phase.
+Colored regions indicate the subspace where the level curves are defined.
+Straight lines correspond to fixed values of the non-normality parameters ($\zeta=\text{const}$ or $\xi=\text{const}$).
+Panel (a) shows the cyclic setup, while panel (b) shows the feedforward setup.
+Increasing $\gamma_{\ell}$ shifts the $m$ level curves leftward, so that $m$ increases for each pair $(\gamma_{\nu},\gamma_{\mu})$, demonstrating that all possible values of $m$ can be obtained while keeping the non-normality parameter fixed.
+Right panels (c–f): Time evolution of $\rho(t)$ under variable reactive conditions.
+Top row (c,d) corresponds to the decoupled system, with couplings fixed and $\xi$ constant.
+Bottom row (e,f) corresponds to setups where the non-normality of feedforward and cyclic systems is adjusted to maintain a consistent level across configurations.
+In this case, only the feedforward system exhibits a significant increase in reactivity.
 Reactivity increases from left to right with the numerical abscissa $m$.
-Here $\zeta$ represents the non-normality parameter for the binary system.
-Top row panels c) and d): In each panel the couplings are chosen for the decoupled system and remain unchanged over the setups.
-From c) and d) $\xi$ keeps fixed.
-Bottom row panels e) and f): As the setup changes, the non-normality of both the feedforward and cyclic systems is adjusted to maintain a consistent level across all configurations in each panel.
-With this change, only the feedforward system exhibits a significant increase in reactivity.
-On the left (panels a) and b)), the level curves of $m$ for different values of $\gamma_{\ell}$ are evaluated at the inactive phase.
-The colored areas correspond to the subspace where the level curves are meaningful.
-The straight lines correspond to the portions of the plane with fixed non-normality ($\zeta=cont$ or $\xi=cont$).
-The cyclic and the feedforward setups are respectively addressed in panels a) and b).
-In both cases, it can be observed that increasing $\gamma_{\ell}$ means shifting the $m$ level curves to the left.
-Then, for each set ($\gamma_{\nu}$,$\gamma_{\mu}$) $m$ is enhanced.
-As a final remark, there it appears clear how all the possible values of $m$ can be obtained while still fixing the non-normality parameter.
 III.3 Reactivity
-Reactivity is one of the distinctive markers of non-normality since it only arises in its presence.
-However, the reverse is not always true, i.e.
-not all non-normal systems exhibit reactivity.
-A reactive system is one in which, during a transient phase the system moves away from the steady state, following a perturbation.
-More specifically, at small time scales the 222-norm of a variable is enhanced before exponentially decaying back to the equilibrium point.
-Formally, reactivity occurs only when the maximum eigenvalue of the Hermitian part of the Jacobian, known as numerical abscissa $m$, is positive:
-$m(\mathbf{J})=\sup\sigma\left[\left(\mathbf{J}+\mathbf{J}^{*}\right)/2\right],$ (19)
-where $\sigma(\mathbf{J})$ denotes the spectrum of the matrix $\mathbf{J}$ and $\mathbf{J}^{*}$ is its conjugate transpose.
-The growth during the transient is proportional to the magnitude of $m(\mathbf{J})$ (as long as $m(\mathbf{J})>0$).
-In the cyclic network, the numerical abscissa can be evaluated analytically. In particular, we find that for the relevant portion of parameter space $\gamma_{\mu},\gamma_{\nu}>0$, we obtain:
+To quantify the immediate transient amplification capability of the network, we calculate the reactivity $\mathcal{R}$, defined as the maximum eigenvalue of the Hermitian part of the Jacobian [12]:
+$\mathcal{R}=\lambda_{\max}\left(\mathcal{H}(J)\right)=\lambda_{\max}\left(\frac{J+J^{\top}}{2}\right).$ (21)
+While the spectral abscissa $s(J)=\max_{i}\text{Re}(\lambda_{i})$ determines the asymptotic stability ($t\to\infty$), the reactivity governs the instantaneous growth rate of the perturbation magnitude at $t=0$.
+This is explicit in the time evolution of the Euclidean norm $\|\delta\mathbf{x}\|^{2}$:
+$\frac{1}{2}\frac{d}{dt}\|\delta\mathbf{x}\|^{2}=\delta\mathbf{x}^{\top}\left(\frac{J+J^{\top}}{2}\right)\delta\mathbf{x}\leq\mathcal{R}\|\delta\mathbf{x}\|^{2}.$ (22)
+A positive reactivity $\mathcal{R}>0$ in an asymptotically stable system ($s(J)<0$) implies a non-monotonic relaxation to the steady state.
+In the language of critical phenomena, this corresponds to a regime of transient fluctuation amplification: even though the system is formally in the disordered (stable) phase, the violation of detailed balance drives a short-time divergence of the susceptibility, allowing fluctuations to grow macroscopically before the asymptotic restoring forces prevail.
+It is crucial to distinguish between non-normality and reactivity.
+While $\mathcal{R}>0$ is a sufficient condition for transient growth, non-normality is a necessary but not sufficient condition for reactivity.
+A system can be non-normal (non-orthogonal eigenvectors) yet non-reactive ($\mathcal{R}<0$), in which case perturbations decay monotonically but potentially with a relaxation time distinct from the asymptotic inverse gap [30, 12].
+Thus, $\mathcal{R}$ specifically measures the contribution of the non-symmetric part of the interactions to the local instability.
+III.3.1 Cyclic couplings
+In this case, the numerical abscissa can be evaluated analytically. In particular, we find that for the relevant portion of parameter space $\gamma_{\mu},\gamma_{\nu}>0$:
 $m=$ $\frac{1}{2}(-2\alpha-2\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}))+$ (23)
 $\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\sqrt{(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu})^{2}+(\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu})^{2}}$
-The results are summarized in Fig.
-3, where we present the level curves of $m$ for several values of $\gamma_{\ell}$ and for $c=0,\epsilon$.
-Notably, panel f) indicates that the behavior of the feed-forward system as a function of $m$ –whose analytic calculation is much more involved– qualitatively matches with that of the cyclic case, shown in panel e).
-This observation suggests that Eq.20 may serve as a qualitative reference even for intermediate configurations with $0<c<\epsilon$.
+Results are summarized in Fig.3(a) and (b), where we present the level curves of $m$ for several values of $\gamma_{\ell}$ and for $c=0,\epsilon$.
+Notably, Fig.3(b) shows the behavior of the feedforward system as a function of $m$, which qualitatively matches that of the cyclic case, as shown in Fig.3(a).
+This observation suggests that Eq.
+(23) may serve as a qualitative reference even for intermediate configurations with $0<c<\epsilon$.
 Interestingly, it turns out that reactivity is governed by two distinct quantities: $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}$ and $\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, with the absolute value of their ratio being the non-normality parameter.
 This implies that, although non-normality is needed to trigger reactivity, a fixed non-normality level does not correspond to a single level of reactivity.
-In other words, the ratio of $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}$ and $\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$ can remain constant while adjusting both numerator and denominator to enhance $m$.
-This is shown explicitly in Fig.
-3 panels e) and f), where the lines representing constant values of $\xi$ or $\zeta$ cross the level curves of $m$, i.e., while moving along a line with constant non-normality, multiple values of reactivity can be achieved.
+In other words, the ratio of $\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu}/\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$ can remain constant while adjusting both numerator and denominator to enhance $m$.
+This is shown explicitly in Fig.3(a) and (b), where the lines representing constant values of $\xi$ or $\zeta$ cross the level curves of $m$, i.e., while moving along a line with constant non-normality, multiple values of reactivity can be achieved.
 Moreover, from Eq.
-20 we define $\Gamma\equiv\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, i.e.
+(23) we can define $\Gamma\equiv\gamma_{\mu}+\epsilon\gamma_{\ell}+\gamma_{\nu}$, i.e.
 the sum of all coupling constants.
-In what follows we fix the non-normality and distinguish between two regimes: one with weak interactions and weak reactivity ($\Gamma,m\ll 1$, illustrated in Fig.
-3 panels c), e)) and one of strong interactions and strong reactivity ($\Gamma,m\gg 1$, illustrated in panels d), f)).
-In the right dashed region, we show the time-evolution of the total activity $\rho(t)$ $=\sqrt{\sum_{i}x_{i}^{2}+y^{2}_{i}}$, starting from an initial condition close to the bisector, where Eq.
-20 holds.
-In the first row, parameters are chosen so that the inner systems are all equivalent, and in particular, the non-normality parameter of the inner systems, $\zeta$, is kept constant.
-We want to compare the evolution of the norm $\rho(t)$ when the reactivity is small (panel c)) with a case in which reactivity is large (panel d)).
-As expected we observe that, when the reactivity is large, the system moves away from the fixed point in 0 before relaxing.
-From these plots, the cyclic system seems to be more reactive than the decoupled system, and that reactivity increases further as $c$ approaches $\epsilon$.
-However, the reactivity may appear larger for the coupled systems, simply because the coupling itself increases the non-normality, which has an effect in the reactivity.
-To isolate the effects of reactivity, in the lower row, we fix the non-normality of each system to the same value.
-In particular we choose parameters in the cyclic system such that $\xi=\zeta$, i.e.
-we adjust $\gamma_{\mu}$ to compensate for the introduction of $\gamma_{\ell}$.
-Again, the left panel, panel e) is for small values of reactivity, and the right panel, panel f) is for larger values of the reactivity, showing larger excursions away from the fixed point.
-Two main messages emerge from these plots: first, we verify that the cyclic system can be reduced to the decoupled case merely through coupling parameter shifts; second, we show that the feedforward system undergoes a genuine increase in reactivity when internal non-normality is held constant.
-For the feedforward system, instead, we do not have a unique global parameter to quantify the total non-normality and summing the parameters for inner and outer non-normality would be an arbitrary choice.
-A specific choice of the couplings that makes the reactivity coincide surely exists, but estimating it is not easy.
-The point that we want to make here is that the external source of non-normality enhances the reactivity.
-In conclusion, the feed-forward network actually increases reactivity compared to the decoupled system.
-In contrast, in the cyclic system the enhancement of reactivity can be fully accounted for by equalizing the level of non-normality.
+In what follows we fix the non-normality and distinguish between two regimes: one with weak interactions and weak reactivity ($\Gamma,m\ll 1$), which is illustrated in Fig.3(c) and (e) and one with strong interactions and strong reactivity ($\Gamma,m\gg 1$), which is illustrated in Fig.3 (d) and (f).
+Fig.3 illustrates the time-evolution of the total activity $\rho(t)$ $\sqrt{\sum_{i}(x_{i}^{2}+y^{2}_{i})}$, starting from an initial condition close to the bisector, where Eq.
+(23) holds.
+In particular, we have chosen parameters allowing the inner systems to be all equivalent, specifically keeping constant the non-normality parameter of the inner systems, $\zeta$.
+Note the comparison of the evolution of the norm $\rho(t)$ when the reactivity is small (see Fig.3(c)), to a case in which reactivity is large (see Fig.3(d)).
+As expected, we observe that for large reactivity, the system moves away from the fixed point with vanishing activity before relaxing.
+From our results, we can conclude that the cyclic system seems to be more reactive than the decoupled system, and that reactivity increases further as $c$ approaches $\epsilon$.
+We must note, however, that the reactivity may appear larger for the coupled systems, simply because the coupling itself increases both the reactivity and the non-normality of the system.
+To isolate the study of the reactivity, we have performed numerical experiments fixing the level of non-normality of the system.
+In particular, we have chosen parameters in the cyclic system such that $\xi=\zeta$, i.e., we fix $\gamma_{\mu}$ to compensate the effects of $\gamma_{\ell}$.
+Therefore, Fig.3(e) reports the results for low values of reactivity, and Fig.3(f) is for high values of reactivity, showing larger excursions away from the fixed point.
+We want to emphasize the two main consequences of our analyses: (i) we have verified that the cyclic system can be reduced to the decoupled case merely through coupling parameter shifts (thus, in the cyclic case, non-reciprocity does not create a new effective dynamical degree of freedom beyond the local E/I module) and (ii) we have shown that the feedforward system undergoes a genuine increase in reactivity when internal non-normality is constant.
+III.3.2 Feedforward couplings
+In contrast, for the feedforward system, there is no unique global parameter to quantify the total non-normality, and simply summing the parameters controlling the inner and outer non-normality would be arbitrary.
+While a specific choice of couplings could, in principle, make the reactivity coincide, identifying it would require careful fine-tuning, outside of the scope of the present work.
+Importantly, the presence of an external non-normal component still enhances the system’s reactive response.
+In conclusion, the feedforward network actually increases reactivity compared to the decoupled system.
+In contrast, in the cyclic system, the enhancement of reactivity can be fully accounted for by matching the level of non-normality.
+Our results show that feedforward architectures provide a robust way to enhance reactivity without increasing local non-normality, a feature that cannot be achieved in symmetric or cyclic networks.
 III.4 Analysis of the stochastic evolution of the dynamics
-As shown in [13, 7], the microscopic underlying system shows very strong finite-size effects.
+As previously shown in [5, 12], the microscopic underlying system shows very strong finite-size effects.
 In particular it was shown in [12] using Eq.
-3, that the demographic noise induces a minimum of the effective potential is the origin, thus generating a bistability in the system.
+(3), that the demographic noise induces a minimum of the effective potential is the origin, thus generating a bistability in the system.
 In presence of strong non-normality, the shear flow along the diagonal, associated with the two nearly-collapsing eigenvectors, facilitates the jumps between the two fixed points and the dynamics resembles the avalanching behaviors observed in neural system, where the global activity undergoes large irregular fluctuations between a low-activity and a high-activity state.
-Here we want to explore the behavior of the coupled system in response to noise. We then couple three mean-field dynamics, where each obeys the system size expansion as in [13]:
+Here, we want to explore the behavior of the coupled system under noise effects. We then couple three E/I populations, where each obeys the system size expansion as in [5]:
 $\frac{dx_{i}}{dt}$ $=-\alpha x_{i}+(1-x_{i})f(s_{i})+\sqrt{\alpha x_{i}+(1-x_{i})f(s_{i})}\eta_{x_{i}}$ (24)
-$\frac{dy_{i}}{dt}$ $=-\alpha y_{i}+(1-y_{i})f(s_{i})+\sqrt{\alpha y_{i}+(1-y_{i})f(s_{i})}\eta_{y_{i}}.$
+$\frac{dy_{i}}{dt}$ $=-\alpha y_{i}+(1-y_{i})f(s_{i})+\sqrt{\alpha y_{i}+(1-y_{i})f(s_{i})}\eta_{y_{i}}.$ (25)
 Also in this case, the introduction of noise induces a new stable equilibrium point in zero for all parameter choices, given that $h$ is kept small.
@@
 4.
-Thanks to the non-normality, similar plots could be obtained for inhibitory activities.
+Similar plots could be obtained for inhibitory activities.
 For low values of the inner non-normality, the system easily gets stuck in its minimum, and excursions between the two minima are not likely.
@@
 When $\gamma_{\ell}<<\gamma_{\mu},\gamma_{\nu}$ the system lies in an incoherent phase (Fig.
-4 panels b) and d)), while when $\gamma_{\ell}\sim\gamma_{\mu},\gamma_{\nu}$ the behavior is more coherent (Fig.
-4 panels a) and c)).
+4 (b) and (d)), while when $\gamma_{\ell}\sim\gamma_{\mu},\gamma_{\nu}$ the behavior is more coherent (Fig.
+4 (a) and (c)).
 When the connectivity among units is large enough and the architecture is cyclic, the activation between units is highly correlated (Fig.
-4a)).
-Instead, in the feed-forward case, when the connectivity among units is large enough, the units are activated sequentially.
+4(a)).
+Instead, in the feedforward case, when the connectivity among units is large enough, the units are activated sequentially.
 To fix the ideas, let us take two adjacent populations, where the bottom one projects input to –but does not receive any input from– the top one.
@@
 The process propagates along the chain, giving rise to the step-like trajectories we observe in Fig.
-4c).
-A discussion about the effective potential induced by the noise and the external couplings is detailed in SM.
+4(c).
+A discussion about the effective potential induced by the noise and the external couplings is detailed in Supp.
+Mat.
+C.
 Figure 4: Comparison of cyclic and feedforward setups in the stochastic regime.
-The top panels illustrate the cyclic configuration, while the bottom ones refer to the feedforward setup.
-The comparison between the left and the right columns shows how the system can be driven from an incoherent regime (on the right, panels b) and d)) to a coherent phase in THE cyclic setup (panel a)) and to step-like trajectories in the feed-forward one (panel c)).
-Parameter choices are as follows: a) $\gamma_{\ell}=1.13$ and b) $\gamma_{\ell}=0.13$; for both $\xi=0.004$.
-c) $\gamma_{\ell}=1.13$ and d) $\gamma_{\ell}=0.07$; for both $\zeta=0.005$.
+Top panels show the cyclic configuration; bottom panels show the feedforward setup.
+Comparing the left and right columns, the system shifts from an incoherent regime (b, d) to either a coherent phase in the cyclic case (a) or step-like trajectories in the feedforward case (c).
+Parameters: (a) $\gamma_{\ell}=1.13$, $\xi=0.004$; (b) $\gamma_{\ell}=0.13$, $\xi=0.004$; (c) $\gamma_{\ell}=1.13$, $\zeta=0.005$; (d) $\gamma_{\ell}=0.07$, $\zeta=0.005$.
 IV Discussion and conclusions
-Firstly, through the analysis of bifurcations and reactivity, we find that systems with feedback are more controllable, meaning that changes in parameter settings – or changes in physiological conditions – result in smoother alterations of network responses compared to purely feedforward systems.
-This enhanced controllability implies that the activation patterns of neurons –which are the substrate for neural representation of the stimulus– are more robust to perturbations when feedback is present.
-Similarly, it has been proposed that when stimuli become consciously reportable, they elicit an ‘ignition process’, which causes information about a brief stimulus to become sustained and broadcasted across many brain areas [30].
-Our work suggests, in an extremely simplified setup, that the presence of feedback might be a key ingredient in such process.
-More specifically, our model reveals that the introduction of feedback into the system plays a critical role in generating coherence between processing stages.
-This observation aligns with experimental findings of ‘zero-lag interaction’ between layers during perception and visual working memory [31, 32].
-These are frequently attributed to either common external inputs or specific network motifs [31, 33].
-Interestingly, our model suggests that feedback alone might suffice to produce this synchrony, even in the absence of external modulation.
-Exploring the influence of time-lagged feedback would be a valuable future direction to test the validity of this hypothesis.
-Additionally, our model shows that in a purely feedforward setup, the stochastic dynamics exhibits a greater degree of variability and complexity in the system’s response patterns, compared to the cyclic setup.
-In the absence of feedback, the neural activity is more distributed across a broader range of activation patterns, reflecting a more diverse set of responses to sensory inputs.
-Remarkably, when the inter-unit connectivity is weak, coherent representations are completely lost, both for feedforward and cyclic networks.
-A priori, we chose $n=3$ to investigate the potential role of complex eigenvalues and to explore whether the system could sustain quasi-cycles under the influence of noise.
-However, our analysis reveals that, despite the theoretical presence of complex eigenvalues, no oscillatory behavior emerges in practice.
-As $n$ increases, the frequency of the oscillation decreases, and the quasi-cycle dynamics potentially predicted by the mathematical formalism might be observable for larger $n$.
-This could be investigated with numerical simulations of systems with $n>3$, but is beyond the scope of the present work.
-This study contributes to the growing body of literature suggesting that feedback plays a pivotal role in the neural dynamics underlying perception.
-Our model provides new insights into how feedback mechanisms could inherently promote synchrony and robustness in neural networks, even in the absence of common input.
+A central motivation of this work, laid out in the Introduction, is that two conceptually distinct sources of non-reciprocity can shape neural dynamics: local dynamical non-normality within each E/I module, and structural non-normality emerging from directed inter-module connectivity.
+Our results clarify when these two ingredients merely “stack” in an effectively reducible way, and when their combination produces genuinely new transient and stochastic behaviors.
+At the local level, each Wilson-Cowan E/I population is tuned to a strongly non-normal regime (captured by the near-alignment of eigenvectors), so that even linearly stable fixed points can display pronounced transient amplification and strong sensitivity to fluctuations.
+This “inner” mechanism, quantified by the parameter $\zeta$, sets the baseline susceptibility of each unit to perturbations and noise-driven excursions between inactive and active states.
+The motif or network level (structural) contribution depends crucially on architecture.
+In the cyclic (feedback) motif, we find that the effect of coupling can be largely reabsorbed as an effective shift of the excitatory coupling (i.e., a renormalization of parameters), so that the deterministic transition and the core local non-normal mechanism are essentially those of the decoupled system in disguise.
+In the language of the Introduction, this is a case where directed interactions do not create an independent “structural” non-normal mechanism that fundamentally reshapes the dynamics; instead, the dominant effect is to move the operating point of the same underlying local non-normal module.
+By contrast, in the purely feedforward motif, directionality cannot be absorbed into a single effective parameter: the network introduces an additional, explicitly “outer” non-normal component (controlled by $\gamma_{\ell}$) on top of the inner E/I non-normality $\zeta$.
+This distinction becomes especially sharp in the analysis of reactivity.
+While non-normality is a prerequisite for reactivity, we show that matching non-normality levels does not uniquely fix reactive amplification; importantly, only the feedforward architecture exhibits a genuine increase in reactivity when inner non-normality is held constant, consistent with the idea that structural non-normality can add an independent amplification channel.
+In other words, this is precisely the kind of “interplay across scales” anticipated in the Introduction: embedding locally non-normal subsystems into a directed architecture can enhance transient responses in a way that is not captured by local properties alone.
+These architectural differences are even more pronounced in the stochastic regime, where demographic noise can generate bistability and, with sufficiently strong non-normality, intermittent avalanche-like excursions.
+Here, structure governs not only the amount of amplification, but also how activity is distributed in time across populations.
+With strong coupling and feedback (the cyclic motif), noise-driven transitions are typically coherent, producing synchronized activation.
+In the feedforward motif, strong coupling instead yields sequential, step-like propagation along the chain, reflecting directional recruitment rather than global coherence.
+Overall, our results distinguish two roles of non-reciprocity: local non-normality permits large noise-driven excursions, while structural asymmetry determines whether those excursions spread synchronously (feedback) or progressively (feedforward).
+Our results also connect to classic work on large random interaction matrices in ecology and neuroscience [26, 33, 7].
+In those settings, departing from symmetric (or nearly symmetric) couplings tends to increase eigenvector non-orthogonality and boost transient amplification, making the dynamics much more sensitive to small perturbations.
+Intuitively, directed (non-reciprocal) interactions can route activity along preferred directions: growth in one subspace can feed into others before relaxation sets in, sustaining irregular fluctuations and, in high dimensions, helping trigger chaotic or strongly time-dependent regimes.
+From this viewpoint, the “outer” non-normality introduced by directed architecture plays a role analogous to asymmetry in random connectivity: it creates extra pathways for redistributing locally amplified perturbations, hindering simple equilibration and promoting persistent, noise-like dynamics (see also [24, 25]).
+Finally, we note limitations and natural extensions.
+We chose $n=3$ to allow complex eigenvalues and to probe noise-driven time dependence, but we did not observe sustained oscillations; larger $n$ may be needed to obtain more robust quasi-cyclic dynamics.
+More broadly, an important next step is to move beyond minimal motifs to richer architectures with many more nodes and heterogeneous connectivity, including mixtures of feedforward pathways, recurrent loops, and multiscale feedback, where local and structural non-normal effects could interact in new ways.
+Testing time-delayed feedback and relaxing the assumption of purely excitatory long-range coupling would further clarify how broadly the separation between inner (local) and outer (structural) non-normality holds in more realistic, large-scale circuits.
+Overall, the present framework supports the Introduction’s central claim: non-reciprocity is layered, and understanding neural dynamics requires disentangling when asymmetry acts through local non-normal mechanisms, when it acts through structural non-normal mechanisms, and when their combination yields new transient and noise-driven collective phenomena.
+We hope this work will encourage further research along these conceptually rich and challenging directions.
+Acknowledgments: We acknowledge the Spanish Ministry and Agencia Estatal de Investigación (AEI), MICIN/AEI/10.13039/501100011033, for financial support, Project PID2023-149174NB-I00 funded also by ERDF/EU.